A Bibliography of Publications in the International Journal of Computer and Information Sciences

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

28 February 2002 Version 1.03

Title word cross-reference

2 [Pro79, ZB81, Zak83]. 3 [ZB81, Zak83]. B [Vou84]. d [DS85]. k [DS85]. m [Iba78]. N [TL85]. ω [JT83a].

-languages [JT83a]. -trees [Pro79].

0L [Roz73]. 0L-systems [Roz73].

1 [Bus72, SKB77, WM72b].

2 [IA72]. 2-SLIDS [IA72].

360 [BMS74]. 360/75 [BMS74].

68 [CTW74].

75 [BMS74].

abduction [Bro82]. abstract [Ell78, Vel82]. acceptor [ET79]. acceptors [San74]. accounting [Sch79]. Ada [Bis85]. Adaptive [LR80, Ang75, Kam75]. adders [YM84]. addition [Tha79]. addressed [Lip72]. addressing [Bis85]. AE1 [Hon85]. aided [BT85, IH74, SI81]. AL [WM72b]. AL-1 [WM72b]. algebra [HK77, Muz77, Sta84, Sto79]. algebraic [Can76, Ras82a, Raś82b]. ALGOL [CTW74, TS72]. ALGOL-68 [CTW74]. algorithm [DI80, ES85, Er84, GET76, Joh76, Lee82, Lee84, MTC77, OT73, Sha81, SD79, SdV77]. Algorithmic [AFC79]. Algorithms [Lin74, YM80, YH84, AIR83, Hug72, Iba76, Iba78, IH74, LS80, LR80, SG72, TS72]. allocation [CL82, Dut85]. alphanumeric [DB78]. Analysis [Ell78, BP83, BH81, Hof78, Isc77, IAT85a, IAT85b, Kam75, LF82, MJ75, TP76]. analyzers [Mis72]. analyzing [DB78]. answer [Min77]. answering [MMF74]. apical [Sub85a, Sub85b]. apical-growth [Sub85a, Sub85b]. APL [Kat72]. applicability [Wey79]. Application [Bra83b, Bra83a, Kan74, Bab73, Dd77, EC79, Jai73, Pay73, Sta76, TF79]. Applications [Bow76, HK77, AR83, SM76]. applied [Iva73, MS74]. approach [AFC79, Bis85, BT85, HK85, Kat72, Lee83b, Ras82a, SS78, Tou80a, Tow80, Raś82b, YH84]. Approximate [Yag81, Hon85]. Approximation [BP83]. arbitrarily [Ros75]. arbitrary [BR81, Ros75]. Architectural [Bis85]. architecture [GG82]. archivable [Kos84]. arrays [Ros75]. assertions [BL81]. assignment [YW75a]. associated [Hug72, Jor77]. associations [Bra83b, Bra83a]. Associative [Wil72, WM72b]. attribute [Boc79, Bra83b, Bra83a, BR73, Nie84]. attribute-oriented [BR73]. attribute-relation [Bra83b, Bra83a]. auditing [HM82]. authorization [Min78]. automata [Dal75a, Dal75b, FK84, JS85a, JS85b, Lee83b, TA77, YM80, YH84, Yeh75]. automated [Jai73]. Automatic [Bie78, TG72, BD72, BD73, Sto79]. automaton [Hur84]. auxiliary [RG75]. availability [TP77]. AVL [Zak83].

backtracking [Sta76]. balance [PIB^{+}75]. balance-of-variables [PIB^{+}75]. balancing [Kam75]. base [AS81a, EJM73, Hin76, KC79, Moy77, WG72, dMF81]. based [AS81a, BS73, GF79, Hoo85, Jor77, Kob81, Min77, Mis72, Nie84, OT73, Tow80, YM84]. bases [AS81b, Nie83]. basic [TL85]. Baxter [Har80]. BCNF [JNS83a, JNS83b]. BD [DS85]. behavior [Jap81, Shn76, SM79, SKB77]. between [Mae76, Mar72]. Binary [ORSW84, SM76, Ste80]. biological [SM76]. bipartite [Gab76, Har80]. bit [AS77, Wan75]. Bitmaps [BR73]. block [BS78, BCM^{+}78a, BCM^{+}78b, BH81, DNR81]. block-structured [BS78, BCM^{+}78a, BCM^{+}78b, DNR81]. Boolean [TP76]. both [DL76]. bound [BD72, BD73, Iba76, Iba78, LMN74]. boundaries [Lan84]. boundary [Tou79a]. bounded [Sal78]. boxes [O’R85]. BPL [Hoo85]. branch [Iba76, Iba78, LMN74]. branch-and-bound [Iba76, Iba78, LMN74]. British [PIB^{+}75]. Bron [Joh76]. Bron-Kerbosch [Joh76]. bubble [Zak81]. buffer [SKB77].

calculi [Kob81]. calculus [Sto79]. canonical [Can79]. capabilities [IA72, Sch84]. capability [WM72a]. causality [FC73]. Cellular [AR83, ET79, Lip72]. CFPG [TF82]. CH [KH84]. chain [MK81, XT82]. chain-like [MK81]. characteristics [DL76, TA77]. Characterizations [Vel80, Sal78]. Characterizing [Fur76]. characters [DB78, TG72]. check [LMN74]. Checking [Ata84]. CHITRA [DB78]. choice [Sto74]. circles [Tou83]. circuits [HK77, SS78]. Circular [Las76]. class [HS85, Kod79, TF82]. classes [Bro82, Dor76, ELR75, OW82a, OW82b, SdV77]. classification [Mae80, PF75, Yag81]. clique [OT73]. clique-detection [OT73]. Cliques [Joh76]. closure [Lee82]. cluster [BWE80, Tou79b]. cluster-seeking [Tou79b]. clustering [Did73, MK81, NH74]. clusters [Did73, MK81]. code [CD84, DI80, Hun72]. codes [Las76, Vel80, Vou82, Vou84]. coexistence [FC73]. cognitive [BCS72, DB78]. collection [BS78]. color [Gab76]. combinational [HK77, Kan74]. combinatory [Sta84]. combined [HR83, Sto74]. communication [Moy77, SKB77, Tre75]. Compacting [Tin77]. compaction [Lan84]. comparative [Zak83]. Comparing [RV82]. comparison [DS85, OW82a, OW82b, ORSW84]. comparisons [Iba76]. compiler [Bus72]. complete [Har80]. complex [Iva73]. complexities [ELR75, Hwa75]. Complexity [CCDK80, Tou84, Iva79, MS85b, San84]. composite [DI80, Har80]. Composition [Ver79]. computability [CCDK80]. computational [Edm77]. Computationally [AIR83]. Computer [Bou74a, Bou74b, Bou77, Duj80, Mis72, SI81, Sto79, Bow76, BS73, BT85, Ell78, FF76, MS74, Min78, Moy75, SKB77, Tou80a, Tre75]. computer-aided [BT85]. Computer-based [Mis72, BS73]. computers [XT82]. Computing [Tou83, APG85, Kam75]. Concatenated [CD84]. concentric [MK81]. Concepts [GG82]. Conceptual [dMF81]. concerning [Vou82]. concurrent [DL76]. condensation [Lin74]. conditional [IAT85a, IAT85b]. configurations [BR81, Tou80a]. confirmation [CLYY85]. connectivity [Fur76]. consecutive [AFC79]. considerations [YW75b]. constraints [Tou83]. constructing [LS80, Ste80]. construction [JS85a, JS85b]. content [Lip72]. content-addressed [Lip72]. context [Boo73, Can76, LF84a, LF84b]. context-free [LF84a, LF84b]. context-sensitive [Boo73, Can76]. contiguous [TP77]. control [Min78, PIB^{+}75, SS83, Sto79]. controlled [Jap81, Sub85a, Sub85b]. controls [HM83]. conversational [SM76]. Conversion [LF82]. convex [Lee83a]. convexity [Tou84]. coordination [Mae80]. copy [Dut85]. coroutines [Sta76]. correcting [FF76]. correctness [Che80]. Cost [BH81, ORSW84, YW75b]. costs [Dut85, ZB81]. count [BCM^{+}78a, BCM^{+}78b]. counting [HC77]. covering [Boc79, HK85, Hon85, IH74, YM85]. coverings [CP84]. criteria [Duj80]. criterion [PIB^{+}75]. CUCH [BD72, BD73]. CUCH-machine [BD72, BD73]. Cut [JMW81a, JMW81b]. cycles [Pro79, PS81]. cyclic [Er84].

dag [Har80]. Data [EJM73, GF79, Lip72, AS80, AS81b, AS81a, BCS72, Cas83, Fur76, GG82, Hin76, Hoa75, Hun72, IA72, Kak83, Kat72, ML80, MS85b, Moy77, Nie83, Nie84, Orm84, RV82, RM85, SS76, Vel82, Wan75, WM72a, WG72, Wil72, dMF81, van72]. database [AS80, CL82, GP83, SI81, Sch84, Zak81]. databases [BBDS80, Dut85, Kos83, Kos84]. deadlock [ES85]. debate [Tre75]. decision [HM82, MJ75, TG72]. decoder [FF76]. decoding [SKB77, VD74]. decomposing [Mae76]. decomposition [Kos84]. deductive [Sch84]. defined [Gut76, OW82a, OW82b]. definitions [Che80]. Degrees [Hug72, TA77]. Delaunay [LS80, Sha81]. dependencies [AS81a, Bra83b, Bra83a, Cas83, GP83]. dependency [AS81b]. dependent [IA72]. deployment [Das77]. Depth [Iba78]. Depth- [Iba78]. derivations [AH79]. described [CTW74]. describing [RM85]. descriptive [Mae80]. design [Bis85, CLM80, Hur84, Koc77, Lei81, NT80, SI81]. designing [LMN74, SS84]. Detection [Mae76, ES85, HK77, Kai77, Kan74, OT73]. Determination [AS77, Tou79a]. determine [Mar72]. deterministic [CW79, ELR75]. Development [Iva79, Whi76]. developmental [ELR75]. diagnosis [Bou74a, Bou74b, Bou77, IAT85a, IAT85b]. diagrams [JNS83a, JNS83b]. diameter [FHP81]. dictionary [Bus72]. difference [TP76]. dimensional [BR81]. dimensions [Ros75]. directed [GF79, MT79, YM80]. distance [Bab73]. distributed [APG85, CL82, Dut85, ES85, Hoo85, San84]. distribution [CM85, Kak84]. distributional [Bro82]. distributions [Dor76]. document [BT85]. documenting [RM85]. domain [BBDS80]. domain-oriented [BBDS80]. domains [Spi73]. dominating [Pro79, PS81]. domination [FHP81]. Doubly [CW79]. DTOL [JMW81a, JMW81b]. dynamic [Did73, RG73, Tou79b]. DYNOC [Tou79b].

ECG [Jai73]. economy [PIB^{+}75]. edge [Gab76, Kai77]. EDP [HM83, HM82]. effective [Gui81]. efficient [AIR83, DNR81, Lee82, MH82, WG72]. elementary [EC79]. elements [TL85]. eliminating [NH74]. Embedding [Sch84]. empirical [DS85]. employing [DL76]. empty [Tou83]. enclosing [O’R85]. encoded [SS76]. Encoding [Tha79, BR81, Sto82, XT82]. Engineering [Tou80b, EJM73, Tou85]. English [Moy77]. Entity [JNS83a, JNS83b]. Entity-relationship [JNS83a, JNS83b]. Enumerating [AH79]. enumeration [ILBM72]. environments [Das77, SD79, Yeh75]. equivalence [Boc79, Woo75]. error [BH82, FF76]. error-correcting [FF76]. estimation [DL76]. Euler [Gab76]. Evaluation [Kob81, BMS74, Das77, Duj80, GM82, Hur84, RG75, TS72]. even [CELT78]. evolutionary [MS85a]. examination [BT85]. examples [Kod79]. execute [Mag79c, Mag79a, Mag79b]. Execution [Ham77]. experiment [Tre75]. experimental [SM79]. experiments [Shn76]. Expert [HM82]. Exploratory [Shn76]. Exponentiation [Kak83]. expression [Lee82]. expressions [BTG83a, BTG83b, Cas83, GM82]. extended [Kob81]. extendible [Hua85, Ros75]. extending [Lan84]. Extension [Roz73, Ear74, Hon85]. extensions [SM76]. external [MH82]. extraction [Bab73, Jai73, Min77, SdV77, TG72].

Familial [Orm84]. families [Vel80]. Fast [Har80]. fault [Dal75a, IAT85a, IAT85b, TP76]. FEAST [Das77]. feature [Das77, Jai73, SdV77, TG72]. features [Bab73]. Field [Isc77]. figures [Ata84]. filamentous [Sub85a, Sub85b]. File [CL82, RG75, Dut85, Put80, RG73, Shn74, SG72]. files [Lei81]. filters [BR73]. Finding [KC79, O’R85, Lee82, Lee83a, YM80]. finite [Hur84, Kod79, Lee83b, TA77, YM80, YH84]. firm [Hin76]. fixed [Lew81]. flexible [Jap81]. flow [MS85b]. Formal [Nie83, APG85, Dd77]. Formalization [Che80]. forms [AS81a, RV82]. formulation [Bro83]. FORTRAN [Gui81, Gut76, Whi76]. free [LF84a, LF84b]. freely [Jor77]. Frequency [Kos84]. full [NH74]. function [Hof78]. functional [Bra83b, Bra83a, GP83]. functionalities [AM82]. functionals [BWE80]. functions [CM80, EC79, Gut76, Kod79, NK77, Nie83]. Further [Vou84]. Fuzzy [Dal75a, Dal75b, HK77, Kan74, NK77]. Fuzzy-state [Dal75a, Dal75b].

G [SKB77, YM84]. G-minimum [YM84]. G/M/1 [SKB77]. game [Nau83]. garbage [BS78]. GASP [HR83]. gates [YM84]. Gaussian [SdV77]. general [Hon85, MTC77]. generalization [YH84]. Generalized [Sch79]. generated [PF75]. generating [Hof78]. generating-function [Hof78]. geometric [BR81]. geometrical [Tou80a]. GGC [HR83]. given [CLYY85]. glimpse [Sta84]. goiters [Bou74a, Bou74b, Bou77]. GPSS [HR83]. GPSS-GASP [HR83]. grammar [Che85b, Sal78]. grammars [BH82, Boc79, Boo73, Che85a, Che85b, FH72, GET76, Gre74, LF84a, LF84b, SJ78]. grammatical [Bud83, Tan84]. graph [Bow76, Joh76, Nau83, Sal78, YM80]. graph-variations [Joh76]. Graphical [AS81a, AS81b]. graphics [van72]. graphs [AH79, GF79, OT73, PS81]. growth [Sub85a, Sub85b].

handling [Isc77]. handprinted [DB78]. handwritten [TG72]. Hanoi [Er84]. hardware [Hur84]. harmonic [Tan84]. hash [CD84]. hashing [Hua85]. hazard [HK77]. hazards [Kan74]. HB [ZB81]. HENTRAN [Gui81]. heuristic [Ang75]. heuristic-adaptive [Ang75]. Heuristically [IH74]. heuristics [YW75a]. Hierarchic [BH82]. hierarchical [Nie83]. high [GG82, Gli83, Shr75]. high-level [Gli83, Shr75]. high-performance [GG82]. Higher [ML80, Wal73]. higher-level [Wal73]. Hoare [DW84]. homogeneous [Tin77]. horizontal [DW84]. hull [Lee83a]. hybrid [MK81]. hypothesis [DL76].

IBM [BMS74]. identity [EC79]. II [BCM^{+}78b, Cla73, Mag79b]. ILLIAC [LR80]. illusions [MS74]. image [AR83]. imperfectly [Bab73]. implementation [BH81, DNR81, NT80, Spi73, WG72, van72]. implicants [CM80]. implicit [Bra83b, Bra83a, ILBM72]. improvement [Spi75]. incompletely [CM80]. independent [AS80]. index [YW75b]. indexed [Lei81, SJ78, Shn74]. indexes [Cla73]. indexing [Bie78]. indices [NT80, Sto74]. Inferability [LF84a, LF84b]. Inference [FK84, TF82, Edm77, GET76, Tan84]. infinite [Sto82]. influence [Nau83].

Information

[Koc77, BCS72, BS73, Orł82b, Hin76, Isc77, MS74, Orl82a, Ras82a, SG72, Raś82b]. Inhomogeneous [AK79, smK79]. input [Bou77, DB78, HT81, SM76]. input-trained [DB78]. input/output [HT81]. insertion [Bie78]. insertions [Ste80]. instructions [Bie78]. integer [ILBM72]. integrated [SS78]. integrity [BL81]. interactions [ELR75, SM79]. interactive [MMF74, RW73, Sta76, van72]. internal [SI81]. interpretation [AS81b]. interpreter [CS80]. interrogation [TS72]. intersecting [HC77]. intersection [Lee84]. intersections [Mar72]. inverse [Sch79, YH84]. inversions [Sto74]. inverted [Put80]. involving [HC77]. isomorphisms [YM80]. issues [BH81, van72]. iterative [Er84, OW82a, OW82b]. IV [Bou74a, LR80].

jobstreams [She80]. join [AD81].

Kerbosch [Joh76]. key [CM85, Kak84].

Kleene [Lee82]. Knowledge

[Tou80b, Tou85].

label [Can79]. labeled [Bab73]. Language [Moy80, AS80, BBDS80, CS80, Gli83, Ham77, Hoo85, Jor77, MV74, WM72b, dLD85]. languages [Mag79c, BS78, BCM^{+}78a, BCM^{+}78b, Bro82, Can79, DNR81, ELR75, FH72, JMW81a, JT83a, JMW81b, JT83b, Mag79a, Mag79b, Roz73, Sal78, San74, Shr75, ST75, Tan84, Wal73, Woo75]. large [Cla73, Kos83, Kos84, NT80, Pay73, SS78]. large-scale [SS78]. largest [Tou83]. lattices [AH79]. learned [Jor77]. Learning [DL79, BS73, DL76, MS85a, SD79, TF79]. Lee [Sha81]. Left [ST75]. Left-noncounting [ST75]. level [Cla73, Gli83, Lei81, Shr75, TG72, Wal73]. libraries [Cla73]. like [BTG83a, BTG83b, MK81]. Lindenmayer [Woo75]. line [RW73, Tre75, Whi76]. linear [CCDK80, SdV77, Tan84, Zal73]. linguistic [Bro82, Bro83]. linguistics [Dd77, Edm77, Moy75]. link [IA72]. link-dependent [IA72]. linked [IA72]. LISP [Kod79]. list [Lin74]. LL [Che85a, Che85b]. location [Tou83]. logic [DW84, Kan74]. loops [CELT78]. LR [BH82].

M [SKB77]. machine [MS85a, Par84, BD72, BD73]. machines [Gli83, Hwa75, Lew81, Zal73]. Madcap [WM72a]. Magnetic [Zak81]. making [Ste80]. management [BCM^{+}78a, BCM^{+}78b, Hin76, Kam75, RG73, Zak81]. manipulation [Kat72, RM85, SM76]. many [Ros75]. Markov [JMW81a, Hug72, JMW81b]. mass [YW75a]. match [Hua85]. matching [FHP81]. mathematical [Edm77, Isc77, Iva73]. matrices [AS77, Lee82, Wan75]. matrix [Hon85, Vou82, Vou84]. maximum [FHP81, Sto79]. means [TF82]. measure [MS85b]. measures [Bab73, Ell78]. mechanisms [Gli83]. memories [Lip72]. memory [Jor77, SG72, SS83, TP77, Zak81]. merging [Put80]. message [CLM80, San84]. method [CM85, Did73, Hon85, Isc77, Kak84, LMN74, MJ75, Nie84, NH74, Ste80]. Methodical [Vel82]. methodology [SS84]. Microcode [Lan84]. microprocessors [Mag79c, Mag79a, Mag79b]. microprogram [SS83]. microprograms [DW84]. Minimal [YM85, KC79, Lee84, MH82, O’R85]. minimization [SS83, YM85]. minimizing [Tin77]. Minimum [Pro79, PS81, YM84]. model [AS80, Ear74, GP83, IAT85a, IAT85b, MS74, MJ74, Moy80, Orm84, PIB^{+}75, Shn74, SM79, Spi73]. Modeling [Dut85, Wal73, Iva73, SS84, dMF81]. models [HS78, Iva79, MS85b, RM85, She80, Yag81]. Modification [Dal75b]. modularity [Spi75]. modulo [Kak83]. monitoring [HM83]. monoid [JT83a, JT83b]. monoids [JS85a, JS85b]. movement [DI80]. MRPPS [MMF74]. multi [TG72]. multi-level [TG72]. multiattribute [HC77, Sch79]. Multidimensional [Hua85, MTC77]. multigraphs [Gab76]. multilevel [RG73]. multipath [BP83]. multiple [CELT78, CM80, Dut85]. multiple-output [CM80]. multiplier [HS85]. multipliers [YM84]. multiprocessing [Koc77]. multiprogrammed [HS78, She80]. multiprogramming [Hof78]. multivalued [GP83]. Multivariate [HS78, She80].

N [Bra83b, Bra83a]. natural

[BBDS80, Jor77]. Near [YW75a, SS83]. Near-optimal [YW75a]. neighborhood [SG72]. neighborhoods [OT73]. net [Jor77]. network [Mag79c, AM82, Mag79a, Mag79b]. networks [FF76, LMN74, Tre75]. noisy [Kai77]. noncounting [ST75]. nondeterministic [CS80]. nonhierarchical [Did73]. nonparametric [Das77, SD79]. NOR [YM84]. normal [AS81a, RV82]. normalization [AD81]. note [CM85, Min77, Sha81, Sub85a, Sub85b, Woo75]. Notes [BD73, Can79]. notion [ET79]. nucleus [CTW74]. number [AS77, Kod79].

object [CLYY85]. objects [Mar72]. Observations [Hin76, Bou74a]. odd [CELT78]. odd-even [CELT78]. off [Whi76]. off-line [Whi76]. OL [CW79]. omega [JT83b]. omega-languages [JT83b]. On-line [Tre75, RW73]. One [San74, AIR83, ILBM72, Lei81]. one-level [Lei81]. one-time [AIR83]. One-way [San74]. online [HM83]. OPAL [SD79]. operating [Yeh75]. operation [Min78, TL85]. operation-control [Min78]. operations [Put80, dMF81]. operator [Muz77]. optical [MS74]. Optimal [Bou74a, HK85, BL81, Bou74b, Iva79, LMN74, Lee84, PIB^{+}75, SD79, SdV77, Tou79b, YW75a]. optimization [DI80, Vou82, Yeh75]. optimizing [GM82, Shn74]. order [ML80, Put80]. ordered [NT80, Sto82]. organization [EJM73, Iva79]. organizing [PIB^{+}75]. oriented [BBDS80, BR73, CLM80]. outerplanar [PS81]. output [CM80, HT81, SM76]. overlaid [Cla73]. Overview [dLD85].

packing [MTC77]. pad [AIR83]. paradise [Sta84]. Parallel [Par84, YM84, Hur84, Tha79, TL85]. parallelism [Mae76]. Parameter [Yeh75]. parameters [BWE80, DL76]. Parsers [SJ78]. parsing [Can76, LF82]. Partial [Pay73, Hua85, Sto74]. partial-match [Hua85]. partially [ES85]. Partitioning [FHP81, SD79]. partitions [Gab76]. pathology [Nau83]. paths [KC79]. PATRICIA [Cla73]. PATRICIA-II [Cla73]. Pattern [Dor76, BP83, DL79, LF82, SdV77, Tho82]. patterns [Ang75, Bab73]. performance [BMS74, DS85, Duj80, GG82, Hur84, Moy80]. permutations [Har80]. physical [Bou74a]. pictures [Kai77]. PL [Bus72]. PL/1 [Bus72]. PLA [YM85]. planar [Ata84, BP83]. planning [Bou74a, TF79]. plans [Ver79]. point [JMW81a, HS78, JMW81b, She80]. points [Lew81]. polygon [Lee83a]. polyhedral [Mar72]. polynomial [Kak83]. position [CLYY85]. possibilistic [Yag81]. power [Mae80]. pragmatic [Spi75, Tow80]. Predicting [TP77]. prediction [PIB^{+}75]. Predictive [XT82]. prefix [Muz77, Vel80]. preprocessor [Gui81]. preserving [SG72]. prime [CM80]. principle [Sto79]. probabilistic [Bab73, Bro82, ET79, SS78]. problem [EC79, Er84, Hon85, MV74, YM85, YW75a].

Problems

[Iva73, FC73, HC77, Lee84, San84, Sch79]. procedure [Ang75, Bro82, Mar72, MMF74, MS85a, MK81]. Process [Shr75, HS78, Mae80, She80, TS72]. processes [Isc77, Mae76]. processing [AR83, BL81, Hin76, Hun72, MS74]. productions [Ver79]. prognosis [Jai73]. program [Bie78, ILBM72, Kod79, KC79, Spi75, Tha79, Wey79]. programmed [LF84a, LF84b]. programmer [Shn76, SM79]. Programming [Gut76, CCDK80, Ham77, ILBM72, SM76, dLD85]. programming-language [Ham77]. programs [CLM80, Ver79, Wey79, Whi76]. proof [MMF74]. Properties [MT79, Wal73]. property [AFC79]. proposal [Spi75]. protection [Bud83, Ell78, Gli83]. protective [Spi73]. prototyping [Hoo85]. prover [CTW74, MV74]. proving [Tow80]. puzzles [CM85, Kak84].

quadratic [NH74]. quadrature [LR80]. queries [Hua85, Kob81, Put80]. Query [ZB81]. Querying [BBDS80]. question [MMF74]. queue [SKB77]. queues [Hof78].

random [Yeh75]. real [Koc77, Whi76]. real-time [Koc77, Whi76]. realization [GG82]. reasoning [Yag81]. recognition [DB78, DL79, Dd77, Har80, MV74, Tho82, TG72]. reconfigurable [SS84]. recovering [Dor76]. recovery [BH81]. rectangle [Lee84]. recurrences [Par84]. Recursive [Hoa75, Che80]. recursively [DB78, Gut76]. reduction [Mag79c, DI80, Mag79a, Mag79b]. Redundancy [LMN74]. reference [BCM^{+}78a, BCM^{+}78b]. refutation [MMF74]. register [CELT78]. Regular [BTG83a, BTG83b, Lee82, Sal78, Sub85a, Sub85b, Vel80]. regular-controlled [Sub85a, Sub85b]. Regular-like [BTG83a, BTG83b]. related [Vel80]. relating [van72]. relation [Bra83b, Bra83a, Fur76]. relational [AS81b, AS81a, Cas83, GP83, Kob81, Sch84, Ver79, Zak81]. relations [FC73]. relationship [JNS83a, JNS83b]. relationships [AM82]. Relevance [Moy75]. reliability [Spi75]. Remarks [Vou82, Vou84]. reordering [HT81]. repair [BH82, HK85]. Representation [Orł82b, Orl82a, Hun72, Kat72]. Representations [MV74]. required [BS78, BCM^{+}78a, BCM^{+}78b]. requirement [MH82]. requirements [MTC77]. resolution [Min77, Tow80]. resolution-based [Min77, Tow80]. resource [MTC77]. resources [Kam75]. Response [She80, HS78]. Response-time [She80]. restrictions [Gre74]. restructuring [Nie83, Nie84]. results [SM79, Wey79]. retention [BS78, BCM^{+}78a, BCM^{+}78b, DNR81]. retrieval [AFC79, BCS72, EJM73, FC73, Ras82a, Raś82b, Wan75]. retrieve [SG72]. retrieving [BS73]. revisited [Tou85]. rewriting [Vel82]. robot [Jap81]. robotic [TF79]. root [AS81a, Ste80]. Rough [Paw82]. rules [TF82]. run [Ear74].

Safety [Bud83]. sample [FK84]. scale [SS78]. scanner [Hur84]. scene [Dd77]. SCFG’s [PF75]. Schachter [Sha81]. Scheduling [HM83, BL81]. Schema [RM85, SI81, Wey79]. schemata [GF79, MT79, Tha79]. scheme [AIR83, BH81, Kod79, Min78, NT80, TL85]. schemes [Mae80]. science [Moy75]. search [Iba76, Iba78, NH74, OW82a, OW82b, ORSW84, Ste80]. searches [BR73]. Searching [Sto82, XT82]. secondary [NH74, YW75b]. security [Kak83]. seeking [Tou79b]. segmentation [Ang75]. selection [CD84, Das77, Duj80, YW75b]. self [IAT85a, IAT85b, Iva79, PIB^{+}75, RM85]. self-describing [RM85]. self-diagnosis [IAT85a, IAT85b]. self-documenting [RM85]. self-organization [Iva79]. self-organizing [PIB^{+}75]. Semantic [Boc79, Mis72, SM79, TF82, Tho82]. semantics [Ham77]. semigroups [Zal73]. Sense [Jap81]. Sense-controlled [Jap81]. sensitive [Boo73, Can76]. separate [Mae76]. separation [Kos83]. Sequential [DL76, PF75, VD74, Hwa75, Lei81, Mae76, Zal73]. serializability [BM85]. series [MJ74]. set [GM82, HK85, Hoo85, IH74, IAT85a, IAT85b, KC79]. set-based [Hoo85]. set-covering [HK85, IH74]. sets [FK84, Fur76, HC77, Paw82, Sto82, TA77]. several [SdV77]. shapes [BP83]. shift [CELT78]. shift-register [CELT78]. similarity [Ata84]. simple [Lee83a, SS83, Moy77]. Simple-English [Moy77]. Simplification [NK77]. Simplified [CELT78]. Simulation [HR82, BMS74, She80, Whi76]. simulations [MS74]. single [SdV77]. size [Bou74b]. SLEUTH [DL76]. SLIDS [IA72]. software [Nie84]. solution [HC77, Par84]. solutions [CCDK80, Sch79]. Some [Gre74, Hwa75, van72, DS85, Ell78, SM76]. sort [CELT78]. sorting [AK79, MH82, Pay73, TL85, smK79]. sorts [TL85]. space [Lee83b, Lee84, MH82, Tin77, YH84]. Specification [Nie84, APG85, CLM80, Vel82]. specifications [RV82]. specified [CM80]. spectra [Woo75]. SQL [Bra83b, Bra83a]. SQL/N [Bra83b, Bra83a]. stability [Dal75a]. stage [IA72]. state [Dal75a, Dal75b, Hur84, Lee83b, YH84]. Stateless [Lew81]. statements [Mae76]. static [Kan74]. Statistical [BWE80, Edm77]. statistics [Iva73]. steam [EJM73]. Stochastic [FH72, MT79, Yeh75]. Storage [YW75b, IA72, RG75, Tin77, Wal73, YW75a]. store [RG73]. storing [Ros75]. strategies [HR82, Iba76]. strategy [Kam75]. strength [DI80]. String [Sal78]. strings [PF75]. structural [Hwa75]. structure [BCS72, Boo73, Bus72, EC79, IA72, Lin74, Nau83, WM72a]. structured [BS78, BCM^{+}78a, BCM^{+}78b, DNR81, Gui81]. Structures [RW73, AS81b, Fur76, Hoa75, Kat72, Lip72, Shn74, SS76, Wil72, Zak81, van72]. Structuring [BS73, WG72]. student [Tre75]. study [SKB77, Zak83]. sub [Che85a, Che85b]. sub-LL [Che85a, Che85b]. subsymptomatologies [Bou74b]. Subword [ELR75]. support [Bis85, HM82]. Surjections [CP84]. survey [Tho82]. switching [AM82, CM80, HK77, Kan74, Muz77, NK77]. synchronization [Las76, Shr75]. Syntactic [JS85a, JS85b, SM79, Tho82, FF76, JT83a, JT83b, Kai77, LF82]. Syntactic/semantic [SM79, Tho82]. syntactical [AH79, Har80, VD74]. Syntax [Ear74, BH82, MT79]. syntax-directed [MT79]. synthesis [Bie78, MJ74]. synthesized [Kod79]. system [APG85, BBDS80, BL81, DB78, Gui81, Hin76, HR83, Hoo85, Iva73, Kam75, Koc77, MMF74, SKB77, Whi76, dLD85]. systems [AS80, AS81a, Bow76, Bud83, CCDK80, CW79, Ell78, JMW81a, HM83, HM82, Hun72, HS78, JMW81b, Kan74, Koc77, Min77, Min78, Put80, Ras82a, Sch84, SS84, Sta76, Sub85a, Sub85b, Tha79, Vel82, Raś82b, Woo75, Zak81, Roz73]. systolic [JS85a, HS85, JS85b]. Szilard [Woo75].

table [NH74]. tabled [CW79, Roz73]. tabular [Wil72]. taxiway [HK85]. teacher [DL76, DL79]. technique [Can76, Das77, Dor76, HT81, Kos83, Kos84, LMN74, Pay73, Tou79b, Tou79a]. techniques [EJM73, TP76, Tho82]. technology [HS85]. temporal [Orł82b, FC73, Orl82a]. testing [SS78]. text [Hun72, Tin77]. texture [MJ74, MJ75]. their [AM82, Dal75a, Dor76, RV82]. theorem [CTW74, EC79, MV74, Tow80]. theorem-prover [CTW74]. Theoretical [Iba76]. theory [Bow76, Cas83, Iva79, MJ75, RG73, SS76]. thesaurus [RW73]. Theta [AD81]. Theta-join [AD81]. thresholding [Tou79a]. Time [BS78, BCM^{+}78a, Kos83, MJ74, AIR83, Ang75, BCM^{+}78b, Ear74, FC73, Hwa75, Koc77, Lee84, She80, Whi76]. time-varying [Hwa75]. times [HS78]. tolerance [Dal75a, Dal75b]. tool [BMS74]. topological [Bro83]. topology [AR83]. Towers [Er84]. TPL [dLD85]. Trace [CLYY85]. traces [Ham77]. trained [DB78]. transaction [BL81]. transducers [LF82]. transform [KH84]. translation [MT79, SS76]. transportable [dLD85]. treatment [BD72, BD73]. tree [BTG83a, BTG83b, DS85, FK84, GET76, JS85a, JS85b, LF82, Par84, TP76]. trees [FHP81, FK84, OW82a, OW82b, ORSW84, Pro79, Ste80, ZB81, Zak83]. triangulation [LS80, Sha81]. trivial [JT83a, JT83b]. turbine [EJM73]. Turing [Lew81]. Two [Cla73, LS80, BR81, IA72]. two-dimensional [BR81]. Two-level [Cla73]. two-stage [IA72]. type [GG82, RV82, Vou84]. type- [Vou84]. types [ML80, Vel82]. typology [Bro83].

unambiguous [AS77, Wan75].

understander [Jor77]. understanding [BBDS80, Tou80a]. unfamiliar [DL76]. unified [WM72a]. unimodality [Tou84]. units [HS85]. unsolvability [Hug72]. unsupervised [Bro82, Das77, SD79]. update [Dut85]. use [DW84, Moy80]. Useless [CM80]. Using [Gab76, CELT78, Ear74, Har80, Iva79, PIB^{+}75, Tre75].

V [Bou77]. validity [BWE80]. variable [Bis85]. variables [BD72, PIB^{+}75, BD73]. variations [DS85, Joh76, Sto79]. various [ELR75]. varying [Hwa75]. vector [Pay73, Tha79]. VEDIC [DL79]. verification [DW84, Gli83]. versa [Moy75]. versus [ZB81]. VI [WM72a]. via [RV82, TG72, Vel82]. vice [Moy75]. view [GP83]. VLSI [Hur84, HS85].

W [Gre74]. W-grammars [Gre74]. waveform [Jai73]. way [San74, SS83]. WCRL [AS80]. WFTA [HT81]. which [JNS83a, JNS83b]. whose [JT83a, JT83b, TL85]. workload [Kam75].

zero [ILBM72]. zero-one [ILBM72]. Zoom [Tou79a]. Zoom-thresholding [Tou79a].

References

Arora:1981:TN

[AD81] S. K. Arora and S. R. Dumpala.

Theta-join and normalization. International Journal of Computer and Information Sciences, 10(4):261–281, August 1981. CO-DEN IJCIAH. ISSN 0091-7036.

Al-Fedaghi:1979:AAC

[AFC79] S. S. Al-Fedaghi and Y. H. Chin.

Algorithmic approach to the consecutive retrieval property. International Journal of Computer and Information Sciences, 8(4): 279–301, August 1979. CODEN IJCIAH. ISSN 0091-7036.

Alter:1979:ESG

[AH79] Ronald Alter and Johnson M.

Hart. Enumerating syntactical graphs and lattices of derivations. International Journal of Computer and Information Sciences, 8(4):261–277, August 1979. CO-DEN IJCIAH. ISSN 0091-7036.

Akritas:1983:CEA

[AIR83] A. G. Akritas, S. S. Iyengar, and A. A. Rampuria. Computationally efficient algorithms for a onetime pad scheme. International Journal of Computer and Information Sciences, 12(4):285–316, August 1983. CODEN IJCIAH. ISSN 0091-7036.

Anisimov:1979:IS

[AK79] A. V. Anisimov and D. E. Knuth. Inhomogeneous sorting. International Journal of Computer and Information Sciences, 8(4):255– 260, August 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Alia:1982:SNF

[AM82] G. Alia and E. Martinelli. On switching network functionalities and their relationships. International Journal of Computer and Information Sciences, 11(2):123– 138, April 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Ang:1975:HAP

[Ang75] Beng-Tung Ang. A heuristicadaptive procedure for segmentation of time patterns. International Journal of Computer and Information Sciences, 4(4):329– 348, December 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Agrawal:1985:TFS

[APG85] V. K. Agrawal, L. M. Patnaik, and P. S. Goel. Towards formal specification of a distributed computing system. International Journal of Computer and Information Sciences, 14(5):277–306, October 1985. CODEN IJCIAH. ISSN 0091-7036.

Ankeney:1983:CTA

[AR83] Lawrence A. Ankeney and Gerhard X. Ritter. Cellular topology and its applications in image processing. International Journal of Computer and Information Sciences, 12(6):433–456, December 1983. CODEN IJCIAH. ISSN 0091-7036.

Aalto:1977:DNU

[AS77] S. Aalto and E. Schuegraf. Determination of the number of unambiguous bit matrices. International Journal of Computer and Information Sciences, 6(3):223– 235, September 1977. CODEN IJCIAH. ISSN 0091-7036.

Arora:1980:WDM

[AS80] S. K. Arora and K. C. Smith.

WCRL: a data model independent language for database systems. International Journal of Computer and Information Sciences, 9(4):287–305, August 1980. CODEN IJCIAH. ISSN 0091-7036.

Arora:1981:GNF

[AS81a] S. K. Arora and K. C. Smith.

Graphical normal forms based on root dependencies in relational data base systems. International Journal of Computer and Information Sciences, 10(4):235–259, August 1981. CODEN IJCIAH. ISSN 0091-7036.

Arora:1981:GID

[AS81b] Sudhir K. Arora and K. C. Smith. A graphical interpretation of dependency structures in relational data bases. International Journal of Computer and Information Sciences, 10(3):187– 213, June 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Atallah:1984:CSP

[Ata84] Mikhail J. Atallah. Checking similarity of planar figures. International Journal of Computer and Information Sciences, 13(4): 279–290, August 1984. CODEN IJCIAH. ISSN 0091-7036.

Babu:1973:APD

[Bab73] C. Chitti Babu. On the application of probabilistic distance measures for the extraction of features from imperfectly labeled patterns. International Journal of Computer and Information Sciences, 2(2):103–114, June 1973. CODEN IJCIAH. ISSN 0091-7036.

Bernorio:1980:QDD

[BBDS80] M. Bernorio, M. Bertoni, A. Dabbene, and M. Somalvico. Querying databases with a domainoriented natural language understanding system. International Journal of Computer and Information Sciences, 9(2):141–159, April 1980. CODEN IJCIAH. ISSN 0091-7036.

Berry:1978:TRRa

[BCM^{+}78a] D. M. Berry, L. M. Chirica, D. F. Martin, A. Sorkin, and J. B.

Johnson. Time required for reference count management in retention block-structured languages. I. International Journal of Computer and Information Sciences, 7(1):11–64, March 1978. CODEN IJCIAH. ISSN 0091-7036.

Berry:1978:TRRb

[BCM^{+}78b] D. M. Berry, L. M. Chirica, D. F.

Martin, A. Sorkin, and J. B. Johnston. On the time required for reference count management in retention block-structured languages. II. International Journal of Computer and Information Sciences, 7(2):91–119, June 1978. CODEN IJCIAH. ISSN 0091-7036.

Biss:1972:DSC

[BCS72] K. O. Biss, R. T. Chien, and F. A. Stahl. A data structure for cognitive information retrieval. International Journal of Computer and Information Sciences, 1(1):17–27, March 1972. CODEN IJCIAH. ISSN 0091-7036.

Bohm:1972:CAT

[BD72] C. Bohm and M. Dezani. A CUCH-machine: the automatic treatment of bound variables. International Journal of Computer and Information Sciences, 1(2): 171–191, June 1972. CODEN IJ-CIAH. ISSN 0091-7036.

Bohm:1973:NCA

[BD73] C. Bohm and M. Dezani. Notes on ‘A CUCH-machine: the automatic treatment of bound variables’. International Journal of Computer and Information Sciences, 2(2):157–160, June 1973. CODEN IJCIAH. ISSN 0091-7036.

Bhargava:1981:CAR

[BH81] B. Bhargava and C. Hua. Cost analysis of recovery block scheme and its implementation issues. International Journal of Computer and Information Sciences, 10(6):359–382, December 1981. CODEN IJCIAH. ISSN 0091-7036.

Barnard:1982:HSE

[BH82] David T. Barnard and Richard C. Holt. Hierarchic syntax error repair for LR grammars. International Journal of Computer and Information Sciences, 11(4):231– 258, August 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Biermann:1978:AII

[Bie78] A. W. Biermann. Automatic insertion of indexing instructions in program synthesis. International Journal of Computer and Information Sciences, 7(1):65–90, March 1978. CODEN IJCIAH. ISSN 0091-7036.

Biswas:1985:ASV

[Bis85] P. Biswas. Architectural support for variable addressing in Ada — a design approach. International Journal of Computer and Information Sciences, 14(1):51– 72, February 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Bhargava:1981:OSI

[BL81] B. Bhargava and L. Lilien. On optimal scheduling of integrity assertions in a transaction processing system. International Journal of Computer and Information Sciences, 10(5):315–330, October 1981. CODEN IJCIAH. ISSN 0091-7036.

Brzozowski:1985:S

[BM85] J. A. Brzozowski and S. Muro.

On serializability. International Journal of Computer and Information Sciences, 14(6):387–403, December 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Bowdon:1974:STP

[BMS74] E. K. Bowdon, Sr., S. A. Mamrak, and F. R. Salz. A simulation tool for performance evaluation of the IBM 360/75. International Journal of Computer and Information Sciences, 3(1): 33–57, March 1974. CODEN IJ-CIAH. ISSN 0091-7036.

Bochmann:1979:SEC

[Boc79] Gregor V. Bochmann. Semantic equivalence of covering attribute grammars. International Journal of Computer and Information Sciences, 8(6):523–539, December 1979. CODEN IJCIAH. ISSN 0091-7036.

Book:1973:SCG

[Boo73] R. V. Book. On the structure of context-sensitive grammars. International Journal of Computer and Information Sciences, 2(2): 129–139, June 1973. CODEN IJ-CIAH. ISSN 0091-7036.

Bouckaert:1974:CDGa

[Bou74a] A. Bouckaert. Computer diagnosis of goiters. IV. optimal planning of physical observations. International Journal of Computer and Information Sciences, 3(1): 71–92, March 1974. CODEN IJ-CIAH. ISSN 0091-7036.

Bouckaert:1974:CDGb

[Bou74b] A. Bouckaert. Computer diagnosis of goiters. the optimal size of optimal subsymptomatologies. International Journal of Computer and Information Sciences, 3(4):345–362, December 1974. CODEN IJCIAH. ISSN 0091-7036.

Bouckaert:1977:CDG

[Bou77] A. Bouckaert. Computer diagnosis of goiters. V. the input. International Journal of Computer and Information Sciences, 6(3): 237–261, September 1977. CO-DEN IJCIAH. ISSN 0091-7036.

Bowie:1976:AGT

[Bow76] W. S. Bowie. Applications of graph theory in computer systems. International Journal of Computer and Information Sciences, 5(1):9–31, March 1976. CODEN IJCIAH. ISSN 0091-7036.

Badii:1983:AMP

[BP83] F. Badi’i and B. Peikari. Approximation of multipath planar shapes in pattern analysis. International Journal of Computer and Information Sciences, 12(2): 99–110, April 1983. CODEN IJ-CIAH. ISSN 0091-7036.

Burke:1973:BFA

[BR73] J. M. Burke and J. T. Rickman. Bitmaps and filters for attribute-oriented searches. International Journal of Computer and Information Sciences, 2(3): 187–200, September 1973. CO-DEN IJCIAH. ISSN 0091-7036.

Boyles:1981:EAT

[BR81] Stephanie M. Boyles and Gerhard X. Ritter. The encoding of arbitrary two-dimensional geometric configurations. International Journal of Computer and Information Sciences, 10(1):1– 25, February 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Bradley:1983:ASA

[Bra83a] J. Bradley. Application of SQL/N to the attribute-relation associations implicit in functional dependencies. International Journal of Computer and Information Sciences, 12(2):65– 86, April 1983. CODEN IJCIAH. ISSN 0091-7036.

Bradley:1983:AAR

[Bra83b] James Bradley. Application of SQL/N to the attribute-relation associations implicit in functional dependencies. International Journal of Computer and Information Sciences, 12(2):65– 86, 1983.

Brody:1982:PUA

[Bro82] E. J. Brody. A procedure for the unsupervised abduction of linguistic distributional classes in probabilistic languages. International Journal of Computer and Information Sciences, 11(3):193– 210, June 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Brody:1983:TFL

[Bro83] E. J. Brody. A topological formulation of linguistic typology. International Journal of Computer and Information Sciences, 12(3): 137–150, June 1983. CODEN IJ-CIAH. ISSN 0091-7036.

Brunnstein:1973:SRI

[BS73] K. Brunnstein and J. W. Schmidt. Structuring and retrieving information in computerbased learning. International Journal of Computer and Information Sciences, 2(2):89–101, June 1973. CODEN IJCIAH. ISSN 0091-7036.

Berry:1978:TRG

[BS78] D. M. Berry and A. Sorkin. Time required for garbage collection in retention block-structured languages. International Journal of Computer and Information Sciences, 7(4):361–404, December 1978. CODEN IJCIAH. ISSN 0091-7036.

Brzakovic:1985:ACD

[BT85] D. Brzakovic and J. T. Tou. An approach to computer-aided document examination. International Journal of Computer and Information Sciences, 14(6):365– 385, December 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Barrero:1983:RLT

[BTG83a] A. Barrero, M. G. Thomason, and R. C. Gonzalez. Regular-like tree expressions. International Journal of Computer and Information Sciences, 12(1):1–11, 1983.

Barrero:1983:RTE

[BTG83b] A. Barrero, M. G. Thomason, and R. C. Gonzalez. Regular-like tree expressions. International Journal of Computer and Information Sciences, 12(1):1–121, February 1983. CODEN IJCIAH. ISSN 0091-7036.

Budd:1983:SGP

[Bud83] Timothy A. Budd. Safety in grammatical protection systems. International Journal of Computer and Information Sciences, 12(6):413–431, December 1983. CODEN IJCIAH. ISSN 0091-7036.

Busam:1972:DSP

[Bus72] V. A. Busam. A dictionary structure for a PL/1 compiler. International Journal of Computer and Information Sciences, 1(3): 235–253, September 1972. CO-DEN IJCIAH. ISSN 0091-7036.

Bezdek:1980:SPC

[BWE80] James C. Bezdek, Michael P.

Windham, and Robert Ehrlich. Statistical parameters of cluster validity functionals. International Journal of Computer and Information Sciences, 9(4):323– 336, August 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Cannon:1976:ATC

[Can76] R. L. Cannon, Jr. An algebraic technique for contextsensitive parsing. International Journal of Computer and Information Sciences, 5(3):257–276, September 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Cannon:1979:NCL

[Can79] Robert L. Cannon, Jr. Notes on canonical label languages. International Journal of Computer and Information Sciences, 8(2): 141–148, April 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Casanova:1983:TDD

[Cas83] Marco A. Casanova. A theory of data dependencies over relational expressions. International Journal of Computer and Information Sciences, 12(3):151–191, June 1983. CODEN IJCIAH. ISSN 0091-7036.

Charnes:1980:CCS

[CCDK80] A. Charnes, W. W. Cooper, S. Duffuaa, and M. Kress. Complexity and computability of solutions to linear programming systems. International Journal of Computer and Information Sciences, 9(6):483–506, December 1980. CODEN IJCIAH. ISSN 0091-7036.

Chang:1984:CHC

[CD84] S. K. Chang and J. C. Dorng. Concatenated hash code selection. International Journal of Computer and Information Sciences, 13(5):395–412, October 1984. CODEN IJCIAH. ISSN 0091-7036.

Chen:1978:SOS

[CELT78] T. C. Chen, K. P. Eswaran, V. Y. Lum, and C. Tung. Simplified odd-even sort using multiple shift-register loops. International Journal of Computer and Information Sciences, 7(3):295– 314, September 1978. CODEN IJCIAH. ISSN 0091-7036.

Chen:1980:FCR

[Che80] Tsong Yueh Chen. Formalization of correctness of recursive definitions. International Journal of Computer and Information Sciences, 9(1):55–61, February 1980. CODEN IJCIAH. ISSN 0091-7036.

Chen:1985:GSG

[Che85a] Jin Hua Chen. LL(1) grammars and sub-LL(1) grammars. International Journal of Computer and Information Sciences, 14(4): 225–234, 1985.

Chen:1985:LGS

[Che85b] Jinhua Chen. LL(1) grammar and Sub-LL(1) grammars. International Journal of Computer and Information Sciences, 14(4): 225–234, August 1985. CODEN IJCIAH. ISSN 0091-7036.

Chang:1982:FAD

[CL82] Shi Kuo Chang and An Chi Liu.

File allocation in a distributed database. International Journal of Computer and Information Sciences, 11(5):325–340, October 1982. CODEN IJCIAH. ISSN 0091-7036.

Clark:1973:PTO

[Cla73] J. L. Clark. PATRICIA-II. twolevel overlaid indexes for large libraries. International Journal of Computer and Information Sciences, 2(4):269–292, December 1973. CODEN IJCIAH. ISSN 0091-7036.

Cunha:1980:DSM

[CLM80] P. R. F. Cunha, C. J. Lucena, and T. S. E. Maibaum. On the design and specification of message oriented programs. International Journal of Computer and Information Sciences, 9(3):161– 191, June 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Chang:1985:TCG

[CLYY85] S. K. Chang, B. Luo, C. C. Yang, and M. C. K. Yang. Trace confirmation with given object position. International Journal of Computer and Information Sciences, 14(6):439–453, December 1985. CODEN IJCIAH. ISSN 0091-7036.

Cutler:1980:UPI

[CM80] Robert B. Cutler and Saburo Muroga. Useless prime implicants of incompletely specified multiple-output switching functions. International Journal of Computer and Information Sciences, 9(4):337–350, August 1980. CODEN IJCIAH. ISSN 0091-7036.

Chan:1985:NMP

[CM85] B. Chan and H. Meijer. A note on the method of puzzles for key distribution. International Journal of Computer and Information Sciences, 14(4):221–223, August 1985. CODEN IJCIAH. ISSN 0091-7036.

Chang:1984:SC

[CP84] Lena Chang and Arthur T. Poe.

Surjections and coverings. International Journal of Computer and Information Sciences, 13(6): 509–520, December 1984. CO-DEN IJCIAH. ISSN 0091-7036.

Cohen:1980:INL

[CS80] J. Cohen and R. Stachel. An interpreter for a nondeterministic language. International Journal of Computer and Information Sciences, 9(3):193–217, June 1980. CODEN IJCIAH. ISSN 0091-7036.

Cohen:1974:NTD

[CTW74] J. Cohen, L. Trilling, and P. Wegner. A nucleus of a theoremprover described in ALGOL-68. International Journal of Computer and Information Sciences, 3(1):1–31, March 1974. CODEN IJCIAH. ISSN 0091-7036.

Culik:1979:DDT

[CW79] K. Culik, II and D. Wood.

Doubly deterministic tabled OL systems. International Journal of Computer and Information Sciences, 8(4):335–347, August 1979. CODEN IJCIAH. ISSN 0091-7036.

DalCin:1975:FAT

[Dal75a] M. Dal Cin. Fuzzy-state automata: their stability and fault tolerance. International Journal of Computer and Information Sciences, 4(1):63–80, March 1975. CODEN IJCIAH. ISSN 0091-7036.

DalCin:1975:MTF

[Dal75b] M. Dal Cin. Modification tolerance of fuzzy-state automata. International Journal of Computer and Information Sciences, 4(1): 81–93, March 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Dasarathy:1977:FFE

[Das77] B. V. Dasarathy. FEAST: feature evaluation and selection technique for deployment in unsupervised nonparametric environments. International Journal of Computer and Information Sciences, 6(4):307–315, December 1977. CODEN IJCIAH. ISSN 0091-7036.

Dasarathy:1978:CCH

[DB78] B. V. Dasarathy and K. P. Bharath Kumar. CHITRA: cognitive handprinted input-trained recursively analyzing system for recognition of alphanumeric characters. International Journal of Computer and Information Sciences, 7(3):253–282, September 1978. CODEN IJCIAH. ISSN 0091-7036.

DiasVelasco:1977:AFL

[Dd77] F. R. Dias Velasco and C. de Renna e Souza. An application of formal linguistics to scene recognition. International Journal of Computer and Information Sciences, 6(4):289–306, December 1977. CODEN IJCIAH. ISSN 0091-7036.

Dhamdhere:1980:CAS

[DI80] D. M. Dhamdhere and J. R. Isaac. A composite algorithm for strength reduction and code movement optimization. International Journal of Computer and Information Sciences, 9(3):243– 273, June 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Diday:1973:DCM

[Did73] E. Diday. The dynamic clusters method in nonhierarchical clustering. International Journal of Computer and Information Sciences, 2(1):61–88, March 1973. CODEN IJCIAH. ISSN 0091-7036.

Dasarathy:1976:SLE

[DL76] B. V. Dasarathy and A. L. Lakshminarasimhan. Sequential learning employing unfamiliar teacher hypothesis (SLEUTH) with concurrent estimation of both the parameters and teacher characteristics. International Journal of Computer and Information Sciences, 5(1):1–7, March 1976. CO-DEN IJCIAH. ISSN 0091-7036.

Dasarathy:1979:LUV

[DL79] B. V. Dasarathy and A. L. Lakshminarasimhan. Learning under a VEDIC teacher (pattern recognition). International Journal of Computer and Information Sciences, 8(1):75–88, February 1979. CODEN IJCIAH. ISSN 0091-7036.

deMaine:1985:TPL

[dLD85] P. A. D. de Maine, S. Leong, and C. G. Davis. A transportable programming language (TPL) system. I overview. International Journal of Computer and Information Sciences, 14(3):161–182, June 1985. CODEN IJCIAH. ISSN 0091-7036.

dosSantos:1981:CMD

[dMF81] C. S. dos Santos, T. S. E. Maibaum, and A. L. Furtado. Conceptual modeling of data base operations. International Journal of Computer and Information Sciences, 10(5):299–314, October 1981. CODEN IJCIAH. ISSN 0091-7036.

DiSanto:1981:EIR

[DNR81] M. Di Santo, L. Nigro, and W. Russo. On the efficient implementation of retention blockstructured languages. International Journal of Computer and Information Sciences, 10(1):39– 54, February 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Dorrough:1976:PCT

[Dor76] D. Dorrough. Pattern classes: a technique for recovering their distributions. International Journal of Computer and Information Sciences, 5(2):165–199, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Dandamudi:1985:EPC

[DS85] S. P. Dandamudi and P. G. Sorenson. An empirical performance comparison of some variations of the k-d tree and BD tree. International Journal of Computer and Information Sciences, 14(3):135–159, June 1985. CODEN IJCIAH. ISSN 0091-7036.

Dujmovic:1980:CSC

[Duj80] J. J. Dujmovic. Computer selection and criteria for computer performance evaluation. International Journal of Computer and Information Sciences, 9(6):435– 458, December 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Dutta:1985:MMC

[Dut85] A. Dutta. Modeling of multiple copy update costs for file allocation in distributed databases. International Journal of Computer and Information Sciences, 14(1): 29–34, February 1985. CODEN IJCIAH. ISSN 0091-7036.

Dasgupta:1984:UHL

[DW84] Subrata Dasgupta and Alan Wagner. The use of Hoare logic in the verification of horizontal microprograms. International Journal of Computer and Information Sciences, 13(6):461–490, December 1984. CODEN IJCIAH. ISSN 0091-7036.

Earley:1974:SEU

[Ear74] J. Earley. Syntax extension using a run time model. International Journal of Computer and Information Sciences, 3(3):189– 196, September 1974. CODEN IJCIAH. ISSN 0091-7036.

Epstein:1979:STE

[EC79] H. I. Epstein and B. F. Caviness. A structure theorem for the elementary functions and its application to the identity problem. International Journal of Computer and Information Sciences, 8(1):9–37, February 1979.

CODEN IJCIAH. ISSN 0091-7036.

Edmundson:1977:SIM

[Edm77] H. P. Edmundson. Statistical inference in mathematical and computational linguistics. International Journal of Computer and Information Sciences, 6(2): 95–129, June 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Ealer:1973:DBO

[EJM73] G. C. Ealer, N. F. Jensen, and E. G. Monaco. Data base organization and retrieval techniques for steam turbine engineering. International Journal of Computer and Information Sciences, 2(4):293–309, December 1973. CODEN IJCIAH. ISSN 0091-7036.

Ellis:1978:ASA

[Ell78] Clarence A. Ellis. Analysis of some abstract measures of protection in computer systems. International Journal of Computer and Information Sciences, 7(3): 219–251, September 1978. CO-DEN IJCIAH. ISSN 0091-7036.

Ehrenfeucht:1975:SCV

[ELR75] A. Ehrenfeucht, K. P. Lee, and G. Rozenberg. Subword complexities of various classes of deterministic developmental languages with interactions. International Journal of Computer and Information Sciences, 4(3):219–236, September 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Er:1984:IAC

[Er84] M. C. Er. An iterative algorithm for the cyclic Towers of Hanoi problem. International Journal of Computer and Information Sciences, 13(2):123–129, April 1984. CODEN IJCIAH. ISSN 0091-7036.

Elmagarmid:1985:PDD

[ES85] A. K. Elmagarmid and A. P.

Sheth. A partially distributed deadlock detection algorithm. International Journal of Computer and Information Sciences, 14(5): 307–330, October 1985. CODEN IJCIAH. ISSN 0091-7036.

Edmundson:1979:NPC

[ET79] H. P. Edmundson and Immanuel I. Tung. The notion of a probabilistic cellular acceptor. International Journal of Computer and Information Sciences, 8(3):181–208, June 1979. CO-DEN IJCIAH. ISSN 0091-7036.

Findler:1973:PTR

[FC73] N. V. Findler and D. Chen. On the problems of time, retrieval of temporal relations, causality, and coexistence. International Journal of Computer and Information Sciences, 2(3):161–185, September 1973. CODEN IJCIAH. ISSN 0091-7036.

Fung:1976:ESD

[FF76] L. W. Fung and K. S. Fu.

An error-correcting syntactic decoder for computer networks. International Journal of Computer and Information Sciences, 5(1):

45–58, March 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Fu:1972:SGL

[FH72] K. S. Fu and T. Huang. Stochastic grammars and languages. International Journal of Computer and Information Sciences, 1(2): 135–170, June 1972. CODEN IJ-CIAH. ISSN 0091-7036.

Farley:1981:PTM

[FHP81] Arthur Farley, Stephen Hedetniemi, and Andrzej Proskurowski. Partitioning trees: matching, domination, and maximum diameter. International Journal of Computer and Information Sciences, 10(1):55–61, February 1981. CODEN IJCIAH. ISSN 0091-7036.

Fukuda:1984:ITA

[FK84] H. Fukuda and K. Kamata. Inference of tree automata from sample sets of trees. International Journal of Computer and Information Sciences, 13(3):177–196, June 1984. CODEN IJCIAH. ISSN 0091-7036.

Furtado:1976:CSD

[Fur76] A. L. Furtado. Characterizing sets of data structures by the connectivity relation. International Journal of Computer and Information Sciences, 5(2):89–109, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Gabow:1976:UEP

[Gab76] H. N. Gabow. Using Euler partitions to edge color bipartite multigraphs. International Journal of Computer and Information Sciences, 5(4):345–355, December 1976. CODEN IJCIAH. ISSN 0091-7036.

Gonzalez:1976:AIT

[GET76] R. C. Gonzalez, J. J. Edwards, and M. G. Thomason. An algorithm for the inference of tree grammars. International Journal of Computer and Information Sciences, 5(2):145–164, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Gotlieb:1979:DSB

[GF79] C. C. Gotlieb and A. L. Furtado.

Data schemata based on directed graphs. International Journal of Computer and Information Sciences, 8(1):39–73, February 1979. CODEN IJCIAH. ISSN 0091-7036.

Giloi:1982:CRH

[GG82] W. K. Giloi and R. Gueth.

Concepts and realization of a high-performance data type architecture. International Journal of Computer and Information Sciences, 11(1):25–54, February 1982. CODEN IJCIAH. ISSN 0091-7036.

Gligor:1983:VPM

[Gli83] V. D. Gligor. The verification of the protection mechanisms of high-level language machines. International Journal of Computer and Information Sciences, 12(4): 211–246, August 1983. CODEN IJCIAH. ISSN 0091-7036.

Grant:1982:OES

[GM82] John Grant and Jack Minker. On optimizing the evaluation of a set of expressions. International Journal of Computer and Information Sciences, 11(3):179–191, June 1982. CODEN IJCIAH. ISSN 0091-7036.

Gyssens:1983:AVF

[GP83] M. Gyssens and J. Paredaens. Another view of functional and multivalued dependencies in the relational database model. International Journal of Computer and Information Sciences, 12(4): 247–267, August 1983. CODEN IJCIAH. ISSN 0091-7036.

Greibach:1974:SRW

[Gre74] S. A. Greibach. Some restrictions on W-grammars. International Journal of Computer and Information Sciences, 3(4):289– 327, December 1974. CODEN IJ-CIAH. ISSN 0091-7036.

Guida:1981:EPS

[Gui81] G. Guida. An effective preprocessor for structured FORTRAN: the HENTRAN system. International Journal of Computer and Information Sciences, 10(4):283– 297, August 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Guttmann:1976:PRD

[Gut76] A. J. Guttmann. Programming recursively defined functions in FORTRAN. International Journal of Computer and Information Sciences, 5(2):111–122, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Hamlet:1977:ETP

[Ham77] R. G. Hamlet. Execution traces and programming-language semantics. International Journal of Computer and Information Sciences, 6(4):263–278, December 1977. CODEN IJCIAH. ISSN 0091-7036.

Hart:1980:FRB

[Har80] Johnson M. Hart. Fast recognition of Baxter permutations using syntactical and complete bipartite composite dag’s. International Journal of Computer and Information Sciences, 9(4):307– 321, August 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Hellerman:1977:SMP

[HC77] H. Hellerman and R. Cavallo.

On the solution of multiattribute problems involving counting in intersecting sets. International Journal of Computer and Information Sciences, 6(1):1–7, March 1977. CODEN IJCIAH. ISSN 0091-7036.

Hinomoto:1976:OFI

[Hin76] H. Hinomoto. Observations of a firm’s information processing with a data base management system. International Journal of Computer and Information Sciences, 5(3):277–302, September 1976. CODEN IJCIAH. ISSN 0091-7036.

Hughes:1977:AFA

[HK77] J. S. Hughes and A. Kandel. Applications of fuzzy algebra to hazard detection in combinational switching circuits. International Journal of Computer and Information Sciences, 6(1):71–82, March 1977. CODEN IJCIAH. ISSN 0091-7036.

Harnett:1985:OTR

[HK85] R. M. Harnett and G. C. Kiel. Optimal taxiway repair: a setcovering approach. International Journal of Computer and Information Sciences, 14(6):405–419, December 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Hassen:1982:ESD

[HM82] J. V. Hassen and W. F. Messier, Jr. Expert systems for decision support in EDP auditing. International Journal of Computer and Information Sciences, 11(5): 357–379, October 1982. CODEN IJCIAH. ISSN 0091-7036.

Hansen:1983:SME

[HM83] J. V. Hansen and W. F. Messier, Jr. Scheduling the monitoring of EDP controls in online systems. International Journal of Computer and Information Sciences, 12(1):35–46, February 1983. CO-DEN IJCIAH. ISSN 0091-7036.

Hoare:1975:RDS

[Hoa75] C. A. R. Hoare. Recursive data structures. International Journal of Computer and Information Sciences, 4(2):105–132, June 1975. CODEN IJCIAH. ISSN 0091-7036.

Hofri:1978:GAM

[Hof78] M. Hofri. A generating-function analysis of multiprogramming queues. International Journal of Computer and Information Sciences, 7(2):121–155, June 1978. CODEN IJCIAH. ISSN 0091-7036.

Hong:1985:AEM

[Hon85] Jia Rong Hong. AE1: an extension matrix approximate method for the general covering problem. International Journal of Computer and Information Sciences, 14(6):421–437, December 1985. CODEN IJCIAH. ISSN 0091-7036.

Hooper:1985:BSL

[Hoo85] J. W. Hooper. BPL: a setbased language for distributed system prototyping. International Journal of Computer and Information Sciences, 14(2):83– 105, April 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Hooper:1982:SS

[HR82] J. W. Hooper and K. D. Reilly.

Simulation strategies. International Journal of Computer and Information Sciences, 11(2):101– 122, April 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Hooper:1983:GCG

[HR83] J. W. Hooper and K. D.

Reilly. The GPSS-GASP combined (GGC) system. International Journal of Computer and Information Sciences, 12(2):111– 136, April 1983. CODEN IJ-CIAH. ISSN 0091-7036.

Hunter:1978:MPP

[HS78] D. W. Hunter and G. S. Shedler.

Multivariate point process models for response times in multiprogrammed systems. International Journal of Computer and Information Sciences, 7(2):193– 217, June 1978. CODEN IJ-CIAH. ISSN 0091-7036.

Hurson:1985:CSM

[HS85] A. R. Hurson and B. Shirazi. A class of systolic multiplier units for VLSI technology. International Journal of Computer and Information Sciences, 14(5):261– 275, October 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Huang:1981:NTW

[HT81] C. H. Huang and F. J. Taylor. A new technique for WFTA input/output reordering. International Journal of Computer and Information Sciences, 10(1):27– 37, February 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Huang:1985:MEH

[Hua85] S.-H. S. Huang. Multidimensional extendible hashing for partial-match queries. International Journal of Computer and Information Sciences, 14(2):73– 82, April 1985. CODEN IJCIAH. ISSN 0091-7036.

Hughes:1972:DUA

[Hug72] C. E. Hughes. Degrees of unsolvability associated with Markov algorithms. International Journal of Computer and Information Sciences, 1(4):355–365, December 1972. CODEN IJCIAH. ISSN 0091-7036.

Hunter:1972:DRC

[Hun72] L. W. Hunter. A data representation code for text processing systems. International Journal of Computer and Information Sciences, 1(1):29–42, March 1972. CODEN IJCIAH. ISSN 0091-7036.

Hurson:1984:VDP

[Hur84] A. R. Hurson. A VLSI design for the parallel finite state automaton and its performance evaluation as a hardware scanner. International Journal of Computer and Information Sciences, 13(6): 491–508, December 1984. CO-DEN IJCIAH. ISSN 0091-7036.

Hwang:1975:SSC

[Hwa75] K. Hwang. Some structural complexities of time-varying sequential machines. International Journal of Computer and Information Sciences, 4(3):237–245, September 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Inselberg:1972:TLS

[IA72] A. D. Inselberg and E. B. Altman. 2-SLIDS: a two-stage linked storage structure with link-dependent data capabilities. International Journal of Computer and Information Sciences, 1(2):115–133, June 1972. CO-DEN IJCIAH. ISSN 0091-7036.

Ishida:1985:ASD

[IAT85a] Y. Ishida, N. Adachi, and H. Tokumaru. An analysis of selfdiagnosis model by conditional fault set. International Journal of Computer and Information Sciences, 14(5):243–259, 1985.

Ishida:1985:ASM

[IAT85b] Y. Ishida, N. Adachi, and H. Tokumaru. An analysis of selfdiagnosis model by conditional fault set. International Journal of Computer and Information Sciences, 14(5):243–259, October 1985. CODEN IJCIAH. ISSN 0091-7036.

Ibaraki:1976:TCS

[Iba76] T. Ibaraki. Theoretical comparisons of search strategies in branch-and-bound algorithms. International Journal of Computer and Information Sciences, 5(4):315–344, December 1976. CODEN IJCIAH. ISSN 0091-7036.

Ibaraki:1978:DSB

[Iba78]  Toshihide Ibaraki.  Depth-m

search in branch-and-bound algorithms. International Journal of Computer and Information Sciences, 7(4):315–343, December 1978. CODEN IJCIAH. ISSN 0091-7036.

Ignizio:1974:HAS

[IH74] J. P. Ignizio and R. M. Harnett. Heuristically aided set-covering algorithms. International Journal of Computer and Information Sciences, 3(1):59–70, March 1974. CODEN IJCIAH. ISSN 0091-7036.

Ibaraki:1972:IEP

[ILBM72] T. Ibaraki, T. K. Liu, C. R.

Baugh, and S. Muroga. An implicit enumeration program for zero-one integer programming. International Journal of Computer and Information Sciences, 1(1):75–92, March 1972. CODEN IJCIAH. ISSN 0091-7036.

Iscovici:1977:FAM

[Isc77] S. Iscovici. Field analysis: a mathematical method of analysis for information handling processes. International Journal of Computer and Information Sciences, 6(2):165–178, June 1977. CODEN IJCIAH. ISSN 0091-7036.

Ivakhnenko:1973:PCS

[Iva73] A. G. Ivakhnenko. Problems of complex system modeling and applied mathematical statistics. International Journal of Computer and Information Sciences, 2(1):49–60, March 1973. CODEN IJCIAH. ISSN 0091-7036.

Ivakhnenko:1979:DMO

[Iva79] A. G. Ivakhnenko. Development of models of optimal complexity using self-organization theory. International Journal of Computer and Information Sciences, 8(2):111–127, April 1979. CO-DEN IJCIAH. ISSN 0091-7036.

Jain:1973:EWF

[Jai73] V. K. Jain. ECG waveform feature extraction and its application to automated prognosis. International Journal of Computer and Information Sciences, 2(3): 231–247, September 1973. CO-DEN IJCIAH. ISSN 0091-7036.

Jappinen:1981:SFR

[Jap81] H. Jappinen. Sense-controlled flexible robot behavior. International Journal of Computer and Information Sciences, 10(2):105– 125, April 1981. CODEN IJ-CIAH. ISSN 0091-7036.

H:1981:CPL

[JMW81a] H. Jürgensen, D. E. Matthews, and D. Wood. Cut point languages of Markov DTOL systems. International Journal of Computer and Information Sciences, 10(5):331–340, 1981.

Jurgensen:1981:CPL

[JMW81b] H. Jurgensen, D. E. Matthews, and D. Wood. Cut point languages of Markov DTOL systems. International Journal of Computer and Information Sciences, 10(5):331–340, October 1981. CODEN IJCIAH. ISSN 0091-7036.

Jajodia:1983:EDW

[JNS83a] S. Jajodia, P. A. Ng, and F. N. Springsteel. Entity-relationship diagrams which are in BCNF. International Journal of Computer and Information Sciences, 12(4): 269–283, August 1983. CODEN IJCIAH. ISSN 0091-7036.

Jajodia:1983:ERD

[JNS83b] Sushil Jajodia, Peter A. Ng, and Frederick N. Springsteel. Entityrelationship diagrams which are in BCNF. International Journal of Computer and Information Sciences, 12(4):269–283, 1983.

Johnston:1976:CGB

[Joh76] H. C. Johnston. Cliques of a graph-variations on the Bron-Kerbosch algorithm. International Journal of Computer and Information Sciences, 5(3):209– 238, September 1976. CODEN IJCIAH. ISSN 0091-7036.

Jordan:1977:NLU

[Jor77] S. R. Jordan. A natural language understander based on a freely associated learned memory net. International Journal of Computer and Information Sciences, 6(1):9–25, March 1977. CODEN IJCIAH. ISSN 0091-7036.

H:1985:SMC

[JS85a] H. Jürgensen and A. Salomaa.

Syntactic monoids in the construction of systolic tree automata. International Journal of Computer and Information Sciences, 14(1):35–49, 1985.

Jurgensen:1985:SMC

[JS85b] H. Jurgensen and A. Salomaa.

Syntactic monoids in the construction of systolic tree automata. International Journal of Computer and Information Sciences, 14(1):35–49, February 1985. CODEN IJCIAH. ISSN 0091-7036.

H:1983:LWS

[JT83a] H. Jürgensen and G. Thierrin.

On ω-languages whose syntactic monoid is trivial. International Journal of Computer and Information Sciences, 12(5):359–365, 1983.

Jurgensen:1983:OWS

[JT83b] H. Jurgensen and G. Thierrin. On omega-languages whose syntactic monoid is trivial. International Journal of Computer and Information Sciences, 12(5):359– 365, October 1983. CODEN IJ-CIAH. ISSN 0091-7036.

Kai:1977:SED

[Kai77] Hwang Kai. On syntactic edge detection in noisy pictures. International Journal of Computer and Information Sciences, 6(1): 27–40, March 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Kak:1983:EMP

[Kak83] Subhash C. Kak. Exponentiation modulo a polynomial for data security. International Journal of Computer and Information Sciences, 12(5):337–346, October 1983. CODEN IJCIAH. ISSN 0091-7036.

Kak:1984:MPK

[Kak84] S. C. Kak. On the method of puzzles for key distribution. International Journal of Computer and Information Sciences, 13(2): 103–109, April 1984. CODEN IJ-CIAH. ISSN 0091-7036.

Kameda:1975:AAW

[Kam75] H. Kameda. The analysis of an adaptive workload balancing strategy in computing system resources management. International Journal of Computer and Information Sciences, 4(4):295– 306, December 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Kandel:1974:AFL

[Kan74] A. Kandel. Application of fuzzy logic to the detection of static hazards in combinational switching systems. International Journal of Computer and Information Sciences, 3(2):129–139, June 1974. CODEN IJCIAH. ISSN 0091-7036.

Katzan:1972:AAR

[Kat72] H. Katzan, Jr. An APL approach to the representation and manipulation of data structures. International Journal of Computer and Information Sciences, 1(2): 93–113, June 1972. CODEN IJ-CIAH. ISSN 0091-7036.

Koh:1979:FMS

[KC79] Hikyoo Koh and Henry Y. H.

Chuang. Finding a minimal set of base paths of a program. International Journal of Computer and Information Sciences, 8(6): 473–488, December 1979. CO-DEN IJCIAH. ISSN 0091-7036.

Kou:1984:CT

[KH84] Wei Dong Kou and Zheng Hu.

CH transform. International Journal of Computer and Information Sciences, 13(4):291–307, August 1984. CODEN IJCIAH. ISSN 0091-7036.

Kobayashi:1981:EQB

[Kob81] I. Kobayashi. Evaluation of queries based on the extended relational calculi. International Journal of Computer and Information Sciences, 10(2):63–103, April 1981. CODEN IJCIAH. ISSN 0091-7036.

Koch:1977:ISD

[Koc77] H. S. Koch. Information system design for real-time multiprocessing systems. International Journal of Computer and Information Sciences, 6(3):193–209, September 1977. CODEN IJCIAH. ISSN 0091-7036.

Kodratoff:1979:CFS

[Kod79] Yves Kodratoff. A class of functions synthesized from a finite number of examples and a LISP program scheme. International Journal of Computer and Information Sciences, 8(6):489–521, December 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Kostovetsky:1983:TST

[Kos83] A. Kostovetsky. Time separation technique for large databases. International Journal of Computer and Information Sciences, 12(3): 193–209, June 1983. CODEN IJ-CIAH. ISSN 0091-7036.

Kostovetsky:1984:FDT

[Kos84] A. Kostovetsky. Frequency decomposition technique for large archivable databases. International Journal of Computer and Information Sciences, 13(4):237– 249, August 1984. CODEN IJ-CIAH. ISSN 0091-7036.

Landskov:1984:MCE

[Lan84] D. Landskov. Microcode compaction; extending the boundaries. International Journal of Computer and Information Sciences, 13(1):1–21, February 1984. CODEN IJCIAH. ISSN 0091-7036.

Lassez:1976:CCS

[Las76] J.-L. Lassez. Circular codes and synchronization. International Journal of Computer and Information Sciences, 5(2):201–208, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Lee:1982:EAF

[Lee82] Edward T. Lee. An efficient algorithm for finding Kleene closure of regular expression matrices. International Journal of Computer and Information Sciences, 11(6):409–415, December 1982. CODEN IJCIAH. ISSN 0091-7036.

Lee:1983:FCH

[Lee83a] D. T. Lee. On finding the convex hull of a simple polygon. International Journal of Computer and Information Sciences, 12(2): 87–98, April 1983. CODEN IJ-CIAH. ISSN 0091-7036.

Lee:1983:SSA

[Lee83b] Tony T. Lee. A state space approach to the finite automata. International Journal of Computer and Information Sciences, 12(5): 317–335, October 1983. CODEN IJCIAH. ISSN 0091-7036.

Lee:1984:OTM

[Lee84] D. T. Lee. An optimal time and minimal space algorithm for rectangle intersection problems. International Journal of Computer and Information Sciences, 13(1): 23–32, February 1984. CODEN IJCIAH. ISSN 0091-7036.

Leipala:1981:DOI

[Lei81] T. Leipala. On the design of onelevel indexed sequential files. International Journal of Computer and Information Sciences, 10(3): 177–186, June 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Lewis:1981:STM

[Lew81] F. D. Lewis. Stateless Turing machines and fixed points. International Journal of Computer and Information Sciences, 10(3): 215–218, June 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Lin:1982:CPT

[LF82] Wei Chung Lin and King Sun Fu. Conversion and parsing of tree transducers for syntactic pattern analysis. International Journal of Computer and Information Sciences, 11(6):417–458, December 1982. CODEN IJCIAH. ISSN 0091-7036.

Lu:1984:ICF

[LF84a] H. R. Lu and K. S. Fu. Inferability of context-free programmed grammars. International Journal of Computer and Information Sciences, 13(1):33–58, 1984.

Lu:1984:ICP

[LF84b] H. R. Lu and K. S. Fu. Inferability of context-free programmed grammars. International Journal of Computer and Information Sciences, 13(1):33–58, February 1984. CODEN IJCIAH. ISSN 0091-7036.

Lindstrom:1974:ALS

[Lin74] G. Lindstrom. Algorithms for list structure condensation. International Journal of Computer and Information Sciences, 3(3):197– 216, September 1974. CODEN IJCIAH. ISSN 0091-7036.

Lipovski:1972:DSC

[Lip72] G. J. Lipovski. Data structures in content-addressed cellular memories. International Journal of Computer and Information Sciences, 1(4):329–353, December 1972. CODEN IJCIAH. ISSN 0091-7036.

Lai:1974:RCT

[LMN74] Hung Chi Lai, S. Muroga, and T. Nakagawa. Redundancy check technique for designing optimal networks by branch-and-bound method. International Journal of Computer and Information Sciences, 3(3):251–271, September 1974. CODEN IJCIAH. ISSN 0091-7036.

Lemme:1980:AQA

[LR80] J. M. Lemme and J. R. Rice.

Adaptive quadrature algorithms for the ILLIAC IV. International Journal of Computer and Information Sciences, 9(1):63–72, February 1980. CODEN IJCIAH. ISSN 0091-7036.

Lee:1980:TAC

[LS80] D. T. Lee and B. J. Schachter.

Two algorithms for constructing a Delaunay triangulation. International Journal of Computer and Information Sciences, 9(3): 219–242, June 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Maekawa:1976:DPB

[Mae76] M. Maekawa. Detection of parallelism between statements by decomposing into separate sequential processes. International Journal of Computer and Information Sciences, 5(3):239–255, September 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Maekawa:1980:CPC

[Mae80] M. Maekawa. A classification of process coordination schemes in descriptive power. International Journal of Computer and Information Sciences, 9(5):383– 406, October 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Mago:1979:NMEa

[Mag79a] G. A. Mago. A network of microprocessors to execute reduction languages. I. International Journal of Computer and Information Sciences, 8(5):349–385, October 1979. CODEN IJCIAH. ISSN 0091-7036.

Mago:1979:NMEb

[Mag79b] G. A. Mago. A network of microprocessors to execute reduction languages. II. International Journal of Computer and Information Sciences, 8(6):435–471, December 1979. CODEN IJCIAH. ISSN 0091-7036.

A:1979:NME

[Mag79c] Gyula A. Magó. A network of microprocessors to execute reduction languages. I. International Journal of Computer and Information Sciences, 8(5):349– 385, 1979.

Maruyama:1972:PDI

[Mar72] K. Maruyama. A procedure to determine intersections between polyhedral objects. International Journal of Computer and Information Sciences, 1(3):255–266, September 1972. CODEN IJ-CIAH. ISSN 0091-7036.

Motzkin:1982:EES

[MH82] Dalia Motzkin and Christina L.

Hansen. An efficient external sorting with minimal space requirement. International Journal of Computer and Information Sciences, 11(6):381–396, December 1982. CODEN IJCIAH. ISSN 0091-7036.

Minker:1977:NAE

[Min77] J. Minker. A note on answer extraction in resolution-based systems. International Journal of Computer and Information Sciences, 6(3):179–192, September 1977. CODEN IJCIAH. ISSN 0091-7036.

Minsky:1978:OSA

[Min78] N. Minsky. An operation-control scheme for authorization in computer systems. International Journal of Computer and Information Sciences, 7(2):157–191, June 1978. CODEN IJCIAH. ISSN 0091-7036.

Mishelevich:1972:CSA

[Mis72] D. J. Mishelevich. Computerbased semantic analyzers. International Journal of Computer and Information Sciences, 1(3): 267–286, September 1972. CO-DEN IJCIAH. ISSN 0091-7036.

McCormick:1974:TSM

[MJ74] B. H. McCormick and S. N. Jayaramamurthy. Time series model for texture synthesis. International Journal of Computer and Information Sciences, 3(4): 329–343, December 1974. CO-DEN IJCIAH. ISSN 0091-7036.

McCormick:1975:DTM

[MJ75] B. H. McCormick and S. N. Jayaramamurthy. A decision theory method for the analysis of texture. International Journal of Computer and Information Sciences, 4(1):1–38, March 1975. CODEN IJCIAH. ISSN 0091-7036.

Murty:1981:HCP

[MK81] M. N. Murty and G. Krishna. A hybrid clustering procedure for concentric and chain-like clusters. International Journal of Computer and Information Sciences, 10(6):397–412, December 1981. CODEN IJCIAH. ISSN 0091-7036.

Maibaum:1980:HOD

[ML80] T. S. E. Maibaum and Carlos J. Lucena. Higher order data types.

International Journal of Computer and Information Sciences, 9(1):31–53, February 1980. CO-DEN IJCIAH. ISSN 0091-7036.

Minker:1974:MIR

[MMF74] J. Minker, J. R. McSkimin, and D. H. Fishman. MRPPS — an interactive refutation proof procedure system for question answering. International Journal of Computer and Information Sciences, 3(2):105–122, June 1974. CODEN IJCIAH. ISSN 0091-7036.

Moyne:1975:RCS

[Moy75] J. A. Moyne. Relevance of computer science to linguistics and vice versa. International Journal of Computer and Information Sciences, 4(3):265–279, September 1975. CODEN IJCIAH. ISSN 0091-7036.

Moyne:1977:SDB

[Moy77] J. A. Moyne. Simple-English for data base communication. International Journal of Computer and Information Sciences, 6(4): 327–343, December 1977. CO-DEN IJCIAH. ISSN 0091-7036.

Moyne:1980:LUP

[Moy80] J. A. Moyne. Language use:

a performance model. International Journal of Computer and Information Sciences, 9(6):459– 481, December 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Mateescu:1974:IPM

[MS74] C. A. Mateescu and I. O. Stamatescu. On an information processing model applied to optical illusions and computer simulations. International Journal of Computer and Information Sciences, 3(2):153–166, June 1974. CODEN IJCIAH. ISSN 0091-7036.

Mlodinow:1985:EPM

[MS85a] L. D. Mlodinow and I. O. Stamatescu. An evolutionary procedure for machine learning. International Journal of Computer and Information Sciences, 14(4): 201–219, August 1985. CODEN IJCIAH. ISSN 0091-7036.

Motteler:1985:CMD

[MS85b] Howard E. Motteler and Carl H. Smith. A complexity measure for data flow models. International Journal of Computer and Information Sciences, 14(2):107–122, April 1985. CODEN IJCIAH. ISSN 0091-7036.

Maryanski:1979:PSS

[MT79] Fred J. Maryanski and Michael G. Thomason. Properties of stochastic syntax-directed translation schemata. International Journal of Computer and Information Sciences, 8(2):89–110, April 1979. CODEN IJCIAH. ISSN 0091-7036.

Maruyama:1977:GPA

[MTC77] K. Maruyama, D. T. Tang, and S. K. Chang. A general packing algorithm for multidimensional resource requirements. International Journal of Computer and Information Sciences, 6(2):131– 149, June 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Muzio:1977:POS

[Muz77] J. C. Muzio. A prefix operator for a switching algebra. International Journal of Computer and Information Sciences, 6(2): 151–163, June 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Minker:1974:RLR

[MV74] J. Minker and G. J. VanderBrug.

Representations of the language recognition problem for a theorem prover. International Journal of Computer and Information Sciences, 3(3):217–250, September 1974. CODEN IJCIAH. ISSN 0091-7036.

Nau:1983:GGS

[Nau83] Dana S. Nau. On game graph structure and its influence on pathology. International Journal of Computer and Information Sciences, 12(6):367–383, December 1983. CODEN IJCIAH. ISSN 0091-7036.

Nishihara:1974:FTQ

[NH74] S. Nishihara and H. Hagiwara.

A full table quadratic search method eliminating secondary clustering. International Journal of Computer and Information Sciences, 3(2):123–128, June 1974. CODEN IJCIAH. ISSN 0091-7036.

Niemi:1983:FRF

[Nie83] Timo Niemi. Formal restructuring functions for hierarchical data bases. International Journal of Computer and Information Sciences, 12(6):385–411, December 1983. CODEN IJCIAH. ISSN 0091-7036.

Niemi:1984:SDR

[Nie84] T. Niemi. Specification of data restructuring software based on the attribute method. International Journal of Computer and Information Sciences, 13(6):425– 460, December 1984. CODEN IJ-CIAH. ISSN 0091-7036.

Neff:1977:SFS

[NK77] T. P. Neff and A. Kandel. Simplification of fuzzy switching functions. International Journal of Computer and Information Sciences, 6(1):55–70, March 1977. CODEN IJCIAH. ISSN 0091-7036.

Ng:1980:DIS

[NT80] Peter A. Ng and Lun Tsai. The design and implementation of a scheme for large ordered indices. International Journal of Computer and Information Sciences, 9(5):407–434, October 1980. CO-DEN IJCIAH. ISSN 0091-7036.

ORourke:1985:FME

[O’R85] Joseph O’Rourke. Finding minimal enclosing boxes. International Journal of Computer and Information Sciences, 14(3):183– 199, June 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Orlowska:1982:RTI

[Orl82a] E. Orlowska. Representation of temporal information. International Journal of Computer and Information Sciences, 11(6):397– 408, December 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Ewa:1982:RTI

[Orł82b] Ewa Orłowska. Representation of temporal information. International Journal of Computer and Information Sciences, 11(6):397– 408, 1982.

Orman:1984:FMD

[Orm84] L. Orman. Familial model of data. International Journal of Computer and Information Sciences, 13(3):149–175, June 1984. CODEN IJCIAH. ISSN 0091-7036.

Ottmann:1984:BST

[ORSW84] Thomas Ottmann, Arnold L.

Rosenberg, Hans-Werner Six, and Derick Wood. Binary search trees with binary comparison cost. International Journal of Computer and Information Sciences, 13(2):77–101, April 1984. CODEN IJCIAH. ISSN 0091-7036.

Osteen:1973:CAB

[OT73] R. E. Osteen and J. T. Tou. A clique-detection algorithm based on neighborhoods in graphs. International Journal of Computer and Information Sciences, 2(4): 257–268, December 1973. CO-DEN IJCIAH. ISSN 0091-7036.

Ottman:1982:CID

[OW82a] Th. Ottman and D. Wood. A comparison of iterative and defined classes of search trees. International Journal of Computer and Information Sciences, 11(3): 155–178, 1982.

Ottmann:1982:CID

[OW82b] Th. Ottmann and D. Wood. A comparison of iterative and defined classes of search trees. International Journal of Computer and Information Sciences, 11(3): 155–178, June 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Pargas:1984:PSR

[Par84] Roy P. Pargas. Parallel solution of recurrences on a tree machine. International Journal of Computer and Information Sciences, 13(4):251–277, August 1984. CO-DEN IJCIAH. ISSN 0091-7036.

Pawlak:1982:RS

[Paw82] Zdzisław Pawlak. Rough sets. International Journal of Computer and Information Sciences, 11(5): 341–356, October 1982. CODEN IJCIAH. ISSN 0091-7036.

Payne:1973:PSL

[Pay73] W. H. Payne. Partial sorting: a large vector technique and its application. International Journal of Computer and Information Sciences, 2(2):141–156, June 1973. CODEN IJCIAH. ISSN 0091-7036.

Persoon:1975:SCS

[PF75] E. Persoon and K. S. Fu. Sequential classification of strings generated by SCFG’s. International Journal of Computer and Information Sciences, 4(3):205– 217, September 1975. CODEN IJCIAH. ISSN 0091-7036.

Parks:1975:SMB

[PIB^{+}75] P. C. Parks, A. G. Ivakhnenko, L. M. Boichuk, Svetalsky, and B. K. A self-organizing model of the British economy for control with optimal prediction using the balance-of-variables criterion. International Journal of Computer and Information Sciences, 4(4): 349–379, December 1975. CO-DEN IJCIAH. ISSN 0091-7036.

Proskurowski:1979:MDC

[Pro79] Andrzej Proskurowski. Minimum dominating cycles in 2-trees. International Journal of Computer and Information Sciences, 8(5): 405–417, October 1979. CODEN IJCIAH. ISSN 0091-7036.

Proskurowski:1981:MDC

[PS81] Andrzej Proskurowski and Maciej M. Sysło. Minimum dominating cycles in outerplanar graphs. International Journal of Computer and Information Sciences, 10(2):127–139, April 1981. CO-DEN IJCIAH. ISSN 0091-7036.

Putkonen:1980:OMO

[Put80] Anne Putkonen. The order of merging operations for queries in inverted file systems. International Journal of Computer and Information Sciences, 9(5):351– 369, October 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Ras:1982:AAI

[Ras82a] Z. W. Ras. An algebraic approach to information retrieval systems. International Journal of Computer and Information Sciences, 11(4):275–293, August 1982. CODEN IJCIAH. ISSN 0091-7036.

W:1982:AAI

[Raś82b] Zbigniew W. Raś. An algebraic approach to information retrieval systems. International Journal of Computer and Information Sciences, 11(4):275–293, 1982.

Rose:1973:TDF

[RG73] L. L. Rose and M. H. Gotterer. A theory of dynamic file management in a multilevel store. International Journal of Computer and Information Sciences, 2(4): 249–256, December 1973. CO-DEN IJCIAH. ISSN 0091-7036.

Rose:1975:FEA

[RG75] L. L. Rose and M. H. Gotterer. File evaluation in auxiliary storage. International Journal of Computer and Information Sciences, 4(3):197–204, September 1975. CODEN IJCIAH. ISSN 0091-7036.

Roussopoulos:1985:SMS

[RM85] N. Roussopoulos and L. Mark. Schema manipulation in selfdescribing and self-documenting data models. International Journal of Computer and Information Sciences, 14(1):1–28, February 1985. CODEN IJCIAH. ISSN 0091-7036.

Rosenberg:1975:SAM

[Ros75] A. L. Rosenberg. On storing arbitrarily many extendible arrays of arbitrary dimensions. International Journal of Computer and Information Sciences, 4(2):189– 196, June 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Rozenberg:1973:ETL

[Roz73] G. Rozenberg. Extension of tabled 0L-systems and languages. International Journal of Computer and Information Sciences, 2(4):311–336, December 1973. CODEN IJCIAH. ISSN 0091-7036.

Remy:1982:CDT

[RV82] J.-L. Rémy and P. A. S. Veloso.

Comparing data type specifications via their normal forms. International Journal of Computer and Information Sciences, 11(3): 141–153, June 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Rickman:1973:SIO

[RW73] J. Rickman and W. E. Walden.

Structures for an interactive online thesaurus. International Journal of Computer and Information Sciences, 2(2):115–127, June 1973. CODEN IJCIAH. ISSN 0091-7036.

Salomon:1978:SGG

[Sal78] K. B. Salomon. String and graph grammar characterizations of bounded regular languages. International Journal of Computer and Information Sciences, 7(1): 1–10, March 1978. CODEN IJ-CIAH. ISSN 0091-7036.

Santos:1974:OAL

[San74] E. S. Santos. One-way acceptors and languages. International Journal of Computer and Information Sciences, 3(2):141– 151, June 1974. CODEN IJ-CIAH. ISSN 0091-7036.

Santoro:1984:MCD

[San84] Nicola Santoro. On the message complexity of distributed problems. International Journal of Computer and Information Sciences, 13(3):131–147, June 1984. CODEN IJCIAH. ISSN 0091-7036.

Schinnar:1979:GIS

[Sch79] Arie P. Schinnar. Generalized inverse solutions of multiattribute accounting problems. International Journal of Computer and Information Sciences, 8(2):129– 139, April 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Schwind:1984:EDC

[Sch84] Camilla B. Schwind. Embedding deductive capabilities in relational database systems. International Journal of Computer and Information Sciences, 13(5): 327–338, October 1984. CODEN IJCIAH. ISSN 0091-7036.

Sheela:1979:ONA

[SD79] B. V. Sheela and B. V. Dasarathy. OPAL: a new algorithm for optimal partitioning and learning in nonparametric unsupervised environments. International Journal of Computer and Information Sciences, 8(3): 239–253, June 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Starks:1977:AOS

[SdV77] S. A. Starks, R. J. P. de Figueiredo, and D. L. Van Rooy.

An algorithm for optimal single linear feature extraction from several Gaussian pattern classes. International Journal of Computer and Information Sciences, 6(1):41–54, March 1977. CODEN IJCIAH. ISSN 0091-7036.

Simon:1972:APN

[SG72] J. C. Simon and G. Guiho. On algorithms preserving neighborhood, to file and retrieve information in a memory. International Journal of Computer and Information Sciences, 1(1):3–15, March 1972. CODEN IJCIAH. ISSN 0091-7036.

Shapiro:1981:NLS

[Sha81] Marvin Shapiro. A note on Lee and Schachter’s algorithm for Delaunay triangulation. International Journal of Computer and Information Sciences, 10(6):413– 418, December 1981. CODEN IJ-CIAH. ISSN 0091-7036.

Shedler:1980:RSM

[She80] G. S. Shedler. Response-time simulation of multivariate point process models for multiprogrammed jobstreams. International Journal of Computer and Information Sciences, 9(2):73– 91, April 1980. CODEN IJCIAH. ISSN 0091-7036.

Shneiderman:1974:MOI

[Shn74] B. Shneiderman. A model for optimizing indexed file structures. International Journal of Computer and Information Sciences, 3(1):93–103, March 1974. CO-DEN IJCIAH. ISSN 0091-7036.

Shneiderman:1976:EEP

[Shn76] B. Shneiderman. Exploratory experiments in programmer behavior. International Journal of Computer and Information Sciences, 5(2):123–143, June 1976. CODEN IJCIAH. ISSN 0091-7036.

Shrivastava:1975:PSH

[Shr75] S. K. Shrivastava. Process synchronization in high-level languages. International Journal of Computer and Information Sciences, 4(2):151–169, June 1975. CODEN IJCIAH. ISSN 0091-7036.

Sarda:1981:CAD

[SI81] N. L. Sarda and J. R. Isaac. Computer aided design of database internal schema. International Journal of Computer and Information Sciences, 10(4):219–234, August 1981. CODEN IJCIAH. ISSN 0091-7036.

Sebesta:1978:PIG

[SJ78] Robert W. Sebesta and Neil D. Jones. Parsers for indexed grammars. International Journal of Computer and Information Sciences, 7(4):345–359, December 1978. CODEN IJCIAH. ISSN 0091-7036.

Srivastava:1977:GMQ

[SKB77] H. M. Srivastava, H. B. Kekre, and Y. N. Bapat. G/M/1 queue study of buffer behavior in the decoding system of computer communication. International Journal of Computer and Information Sciences, 6(4):317– 326, December 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Skavaril:1976:BIO

[SM76] R. V. Skavaril and O. R. Miller.

Binary input, output, and manipulation extensions of conversational programming with some biological applications. International Journal of Computer and Information Sciences, 5(4):357– 373, December 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Shneiderman:1979:SSI

[SM79] B. Shneiderman and R. Mayer.

Syntactic/semantic interactions in programmer behavior: a model and experimental results. International Journal of Computer and Information Sciences, 8(3):219–238, June 1979. CO-DEN IJCIAH. ISSN 0091-7036.

V:1979:IS

[smK79] A. V. Anīsīmov and D. E. Knuth.

Inhomogeneous sorting. International Journal of Computer and Information Sciences, 8(4):255– 260, 1979.

Spier:1973:MIP

[Spi73] M. J. Spier. A model implementation for protective domains. International Journal of Computer and Information Sciences, 2(3): 201–229, September 1973. CO-DEN IJCIAH. ISSN 0091-7036.

Spier:1975:PPI

[Spi75] M. J. Spier. A pragmatic proposal for the improvement of program modularity and reliability.

International Journal of Computer and Information Sciences, 4(2):133–149, June 1975. CO-DEN IJCIAH. ISSN 0091-7036.

Shneiderman:1976:TTE

[SS76] B. Shneiderman and S. C. Shapiro. Toward a theory of encoded data structures and data translation. International Journal of Computer and Information Sciences, 5(1):33–43, March 1976. CODEN IJCIAH. ISSN 0091-7036.

Singh:1978:PAT

[SS78] Shanker Singh and Vijendra Pal Singh. A probabilistic approach for testing large-scale integrated circuits. International Journal of Computer and Information Sciences, 7(3):283–294, September 1978. CODEN IJCIAH. ISSN 0091-7036.

Srimani:1983:SWN

[SS83] P. K. Srimani and B. P. Sinha. A simple way to near minimization of microprogram control memory. International Journal of Computer and Information Sciences, 12(1):47–64, February 1983. CO-DEN IJCIAH. ISSN 0091-7036.

Srini:1984:MDM

[SS84] V. P. Srini and B. D. Shrivr. A methodology for designing and modeling reconfigurable systems. International Journal of Computer and Information Sciences, 13(5):339–394, October 1984. CODEN IJCIAH. ISSN 0091-7036.

Shyr:1975:LL

[ST75] H. J. Shyr and G. Thierrin. Leftnoncounting languages. International Journal of Computer and Information Sciences, 4(1):95– 102, March 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Stallings:1976:ACB

[Sta76] W. Stallings. An application of coroutines and backtracking in interactive systems. International Journal of Computer and Information Sciences, 5(4):303– 313, December 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Stark:1984:GPC

[Sta84] W. Richard Stark. A glimpse into the paradise of combinatory algebra. International Journal of Computer and Information Sciences, 13(3):219–236, June 1984. CODEN IJCIAH. ISSN 0091-7036.

Stephenson:1980:MCB

[Ste80] C. J. Stephenson. A method for constructing binary search trees by making insertions at the root. International Journal of Computer and Information Sciences, 9(1):15–29, February 1980. CO-DEN IJCIAH. ISSN 0091-7036.

Stonebraker:1974:CPI

[Sto74] M. Stonebraker. The choice of partial inversions and combined indices. International Journal of Computer and Information Sciences, 3(2):167–188, June 1974. CODEN IJCIAH. ISSN 0091-

Stoutemyer:1979:CAC

[Sto79] D. R. Stoutemyer. Computer algebra for the calculus of variations, the maximum principle and automatic control. International Journal of Computer and Information Sciences, 8(5):419– 434, October 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Stout:1982:SEI

[Sto82] Quentin F. Stout. Searching and encoding for infinite ordered sets. International Journal of Computer and Information Sciences, 11(1):55–72, February 1982. CO-DEN IJCIAH. ISSN 0091-7036.

Subramanian:1985:NRA

[Sub85a] K. G. Subramanian. A note on regular-controlled apical-growth filamentous systems. International Journal of Computer and Information Sciences, 14(4):235– 242, August 1985. CODEN IJ-CIAH. ISSN 0091-7036.

Subramanian:1985:NRC

[Sub85b] K. G. Subramanian. A note on regular-controlled apical-growth filamentous systems. International Journal of Computer and Information Sciences, 14(4):235– 242, 1985.

Tanatsugu:1977:CSD

[TA77] K. Tanatsugu and S. Arikawa. On characteristics sets and degrees of finite automata. International Journal of Computer and Information Sciences, 6(1): 83–93, March 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Tanatsugu:1984:GIH

[Tan84] Keisuke Tanatsugu. A grammatical inference for harmonic linear languages. International Journal of Computer and Information Sciences, 13(5):413–423, October 1984. CODEN IJCIAH. ISSN 0091-7036.

Tangwongsan:1979:ALR

[TF79] S. Tangwongsan and K. S. Fu. An application of learning to robotic planning. International Journal of Computer and Information Sciences, 8(4):303–333, August 1979. CODEN IJCIAH. ISSN 0091-7036.

Tai:1982:ICC

[TF82] J. W. Tai and K. S. Fu. Inference of a class of CFPG by means of semantic rules. International Journal of Computer and Information Sciences, 11(1): 1–23, February 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Tou:1972:ARH

[TG72] J. T. Tou and R. C. Gonzalez.

Automatic recognition of handwritten characters via feature extraction and multi-level decision. International Journal of Computer and Information Sciences, 1(1):43–65, March 1972. CODEN IJCIAH. ISSN 0091-7036.

Thayse:1979:EPP

[Tha79] A. Thayse. Encoding of parallel program schemata by vector addition systems. International Journal of Computer and Information Sciences, 8(3):209– 218, June 1979. CODEN IJ-CIAH. ISSN 0091-7036.

Thomason:1982:SST

[Tho82] M. G. Thomason. Syntactic/ semantic techniques in pattern recognition: a survey. International Journal of Computer and Information Sciences, 11(2):75– 100, April 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Ting:1977:CHT

[Tin77] T. C. Ting. Compacting homogeneous text for minimizing storage space. International Journal of Computer and Information Sciences, 6(3):211–221, September 1977. CODEN IJCIAH. ISSN 0091-7036.

Tseng:1985:PSS

[TL85] S. S. Tseng and R. C. T. Lee. A parallel sorting scheme whose basic operation sorts N elements. International Journal of Computer and Information Sciences, 14(6):455–467, December 1985. CODEN IJCIAH. ISSN 0091-7036.

Tou:1979:ZTB

[Tou79a] J. T. Tou. Zoom-thresholding technique for boundary determination. International Journal of Computer and Information Sciences, 8(1):3–8, February 1979. CODEN IJCIAH. ISSN 0091-7036.

Tou:1979:DDO

[Tou79b] Julius T. Tou. DYNOC—a dynamic optimal cluster-seeking technique. International Journal of Computer and Information Sciences, 8(6):541–547, December 1979. CODEN IJCIAH. ISSN 0091-7036.

Tou:1980:AUG

[Tou80a] J. T. Tou. An approach to understanding geometrical configurations by computer. International Journal of Computer and Information Sciences, 9(1):1–13, February 1980. CODEN IJCIAH. ISSN 0091-7036.

Tou:1980:KE

[Tou80b] J. T. Tou. Knowledge engineering. International Journal of Computer and Information Sciences, 9(4):275–285, August 1980. CODEN IJCIAH. ISSN 0091-7036.

Toussaint:1983:CLE

[Tou83] Godfried T. Toussaint. Computing largest empty circles with location constraints. International Journal of Computer and Information Sciences, 12(5):347–358, October 1983. CODEN IJCIAH. ISSN 0091-7036.

Toussaint:1984:CCU

[Tou84] Godfried T. Toussaint. Complexity, convexity, and unimodality. International Journal of Computer and Information Sciences, 13(3):197–217, June 1984. CO-DEN IJCIAH. ISSN 0091-7036.

Tou:1985:KER

[Tou85] J. T. Tou. Knowledge engineering revisited. International Journal of Computer and Information Sciences, 14(3):123–134, June 1985. CODEN IJCIAH. ISSN 0091-7036.

Townley:1980:PAR

[Tow80] Judy A. Townley. A pragmatic approach to resolution-based theorem proving. International Journal of Computer and Information Sciences, 9(2):93–116, April 1980. CODEN IJCIAH. ISSN 0091-7036.

Thomason:1976:BDT

[TP76] M. G. Thomason and E. W. Page. Boolean difference techniques in fault tree analysis. International Journal of Computer and Information Sciences, 5(1): 81–88, March 1976. CODEN IJ-CIAH. ISSN 0091-7036.

Thesen:1977:PAC

[TP77] A. Thesen and T. Pinkerton. Predicting the availability of contiguous memory. International Journal of Computer and Information Sciences, 6(4):279–287, December 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Treu:1975:OSD

[Tre75] S. Treu. On-line student debate: an experiment in communication using computer networks. International Journal of Computer and Information Sciences, 4(1): 39–51, March 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Terrenoire:1972:EAI

[TS72] M. Terrenoire and M. Simonet. An evaluation of ALGOL 68 for interrogation process algorithms. International Journal of Computer and Information Sciences, 1(1):67–73, March 1972. CODEN IJCIAH. ISSN 0091-7036.

vanDam:1972:SII

[van72] A. van Dam. Some implementation issues relating to data structures for interactive graphics. International Journal of Computer and Information Sciences, 1(4): 287–314, December 1972. CO-DEN IJCIAH. ISSN 0091-7036.

Velasco:1974:SSD

[VD74] F. R. D. Velasco and C. De-Renna e. Souza. Sequential syntactical decoding. International Journal of Computer and Information Sciences, 3(4):273–287, December 1974. CODEN IJ-CIAH. ISSN 0091-7036.

Veloso:1980:CRP

[Vel80] Paulo A. S. Veloso. Characterizations for the regular prefix codes and related families. International Journal of Computer and Information Sciences, 9(5):371– 382, October 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Veloso:1982:MSA

[Vel82] P. A. S. Veloso. Methodical specification of abstract data types via rewriting systems. International Journal of Computer and Information Sciences, 11(5):295–323, October 1982. CODEN IJCIAH. ISSN 0091-7036.

Vere:1979:CRP

[Ver79] Steven A. Vere. Composition of relational productions for plans and programs. International Journal of Computer and Information Sciences, 8(2):149–179, April 1979. CODEN IJCIAH. ISSN 0091-7036.

Voukalis:1982:RCO

[Vou82] D. C. Voukalis. Remarks concerning the optimization of matrix codes. International Journal of Computer and Information Sciences, 11(4):259–273, August 1982. CODEN IJCIAH. ISSN 0091-7036.

Voukalis:1984:FRM

[Vou84] D. C. Voukalis. Further remarks for the matrix type-B codes. International Journal of Computer and Information Sciences, 13(4): 309–326, August 1984. CODEN IJCIAH. ISSN 0091-7036.

Walk:1973:MSP

[Wal73] K. Walk. Modeling of storage properties of higher-level languages. International Journal of Computer and Information Sciences, 2(1):1–24, March 1973. CODEN IJCIAH. ISSN 0091-7036.

Wang:1975:DRU

[Wan75] Y. R. Wang. On data retrieval from unambiguous bit matrices. International Journal of Computer and Information Sciences, 4(2):171–187, June 1975. CO-DEN IJCIAH. ISSN 0091-7036.

Weyuker:1979:APS

[Wey79] Elaine J. Weyuker. The applicability of program schema results to programs. International Journal of Computer and Information Sciences, 8(5):387–403, October 1979. CODEN IJCIAH. ISSN 0091-7036.

Wile:1972:IBE

[WG72] D. S. Wile and C. M. Geschke.

An implementation base for efficient data structuring. International Journal of Computer and Information Sciences, 1(3):209– 224, September 1972. CODEN IJCIAH. ISSN 0091-7036.

White:1976:OSS

[Whi76] J. W. White. An off-line simulation system for development of real-time FORTRAN programs. International Journal of Computer and Information Sciences, 5(1):59–79, March 1976. CODEN IJCIAH. ISSN 0091-7036.

Wilkes:1972:ATD

[Wil72] M. V. Wilkes. Associative tabular data structures. International Journal of Computer and Information Sciences, 1(3):225– 233, September 1972. CODEN IJCIAH. ISSN 0091-7036.

Wells:1972:UDS

[WM72a] M. B. Wells and J. B. Morris.

The unified data structure capability in Madcap VI. International Journal of Computer and Information Sciences, 1(3):193– 208, September 1972. CODEN IJCIAH. ISSN 0091-7036.

Wilkes:1972:ALA

[WM72b] M. V. Wilkes and W. D.

Manville. The associative language AL-1. International Journal of Computer and Information Sciences, 1(4):315–327, December 1972. CODEN IJCIAH. ISSN 0091-7036.

Wood:1975:NLS

[Woo75] D. Wood. A note on Lindenmayer systems, Szilard languages, spectra, and equivalence. International Journal of Computer and Information Sciences, 4(1):53– 62, March 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Xu:1982:PSC

[XT82] Jian Hua Xu and Julius T. Tou. Predictive searching for chain encoding by computers. International Journal of Computer and Information Sciences, 11(4):213– 229, August 1982. CODEN IJ-CIAH. ISSN 0091-7036.

Yager:1981:ARP

[Yag81] Ronald R. Yager. Approximate reasoning and possibilistic models in classification. International Journal of Computer and Information Sciences, 10(2):141–175, April 1981. CODEN IJCIAH. ISSN 0091-7036.

Yeh:1975:POS

[Yeh75] Hsi-Han Yeh. Parameter optimization of stochastic automata operating in random environments. International Journal of Computer and Information Sciences, 4(3):247–263, September 1975. CODEN IJCIAH. ISSN 0091-7036.

Yang:1984:AIG

[YH84] Chao Chih Yang and Hou Kuan Huang. Algorithms for the inverse and a generalization of the state space approach to finite automata. International Journal of Computer and Information Sciences, 13(1):59–76, February 1984. CODEN IJCIAH. ISSN 0091-7036.

Yang:1980:AFD

[YM80] Chao Chih Yang and Charmane P. May. Algorithms for finding directed graph isomorphisms by finite automata. International Journal of Computer and Information Sciences, 9(2): 117–140, April 1980. CODEN IJ-CIAH. ISSN 0091-7036.

Yu:1984:PMN

[YM84] Ging-Shung Yu and S. Muroga.

Parallel multipliers with NOR gates based on G-minimum adders. International Journal of Computer and Information Sciences, 13(2):111–121, April 1984. CODEN IJCIAH. ISSN 0091-7036.

Young:1985:MCP

[YM85] Ming Huei Young and Saburo Muroga. Minimal covering problem and PLA minimization. International Journal of Computer and Information Sciences, 14(6): 337–364, December 1985. CO-DEN IJCIAH. ISSN 0091-7036.

Yue:1975:NHA

[YW75a] P. C. Yue and C. K. Wong.

Near-optimal heuristics for an assignment problem in mass storage. International Journal of Computer and Information Sciences, 4(4):281–294, December 1975. CODEN IJCIAH. ISSN 0091-7036.

Yue:1975:SCC

[YW75b] P. C. Yue and C. K. Wong. Storage cost considerations in secondary index selection. International Journal of Computer and Information Sciences, 4(4):307– 327, December 1975. CODEN IJ-CIAH. ISSN 0091-7036.

Zaki:1981:MBM

[Zak81]  M. Zaki.  Magnetic bubble

memory structures for relational database management systems. International Journal of Computer and Information Sciences, 10(5):341–358, October 1981. CODEN IJCIAH. ISSN 0091-7036.

Zaki:1983:CST

[Zak83] A. S. Zaki. A comparative study of 2-3 trees and AVL trees. International Journal of Computer and Information Sciences, 12(1): 13–33, February 1983. CODEN IJCIAH. ISSN 0091-7036.

Zalcstein:1973:SLS

[Zal73] Y. Zalcstein. On the semigroups of linear sequential machines. International Journal of Computer and Information Sciences, 2(1): 25–28, March 1973. CODEN IJ-CIAH. ISSN 0091-7036.

Zaki:1981:QCH

[ZB81] A. S. Zaki and J. L. Baer. Query costs in HB(1) trees versus 2-3 trees. International Journal of Computer and Information Sciences, 10(6):383–395, December 1981. CODEN IJCIAH. ISSN 0091-7036.